### v1

Abstract
A longstanding open question in algorithm design is whether “combinatorial" matrix multiplication algorithms—avoiding Strassen-like divide-and-conquer—can achieve truly subcubic runtime $n^{3-\delta}$. We present an $O(n^{2.89})$-time exact algorithm, which only sums convolutions in $\mathbb{Z}_{m}^{k}$ (multivariate polynomial multiplications) via FFT, building on the work of Cohn, Kleinberg, Szegedy and Umans (CKSU’05). While the algorithm avoids recursion, the asymptotic speedup arises only for impractically large matrices.
Motivated by practical applications, we use this baseline to develop a new framework for fast approximate matrix multiplication (AMM), via low-degree approximations of the CKSU polynomials. We show that combining the aforementioned algorithm with black-box linear sketching already breaks the longstanding linear speed-accuracy tradeoff for AMM (Sarlos’06, Clarkson-Woodruff’13, Pagh’11, Cohn-Lewis’00), achieving $\frac{1}{r^{1.1}}\|\mathbf{A}\|_{F}^{2}\|\mathbf{B}\|_{F}^{2}$ error in $O(rn^{2})$-time.
Our main result is a low-degree approximation scheme for the CKSU polynomials, based on a Fourier-concentration lemma, yielding substantially smaller error in the distributional setting where $\mathbf{A},\mathbf{B}$ come from an i.i.d product-distribution; For random Gaussian matrices, this practical AMM algorithm attains smaller error than the best rank-$r$ SVD of the output matrix $\mathbf{A}\mathbf{B}$, in time $O(rn^{2})$. This is a substantial improvement over iterative Krylov subspace methods for low-rank approximation. Our theoretical and empirical results suggest the possibility of replacing MatMuls with sums of convolutions in LLM training and inference.
1 Introduction
Matrix multiplication (MatMul) is a fundamental operation across all fields of science and technology, and underlies most industry-scale applications. Above all, MatMuls are the computation and performance bottleneck of training and inference of deep neural networks, where both forward and backpropagation rely on giant matrix multiplications — for example, multiplying $16K\times 16K$ matrices is now considered a prerequisite in any LLM [al24, Var23]. As such, the continual increase in computations and energy facilitating AI breakthroughs poses a real problem of scalability for the field.
The discovery of fast matrix multiplication (FMM) [Str86, Wil12, CW13, DWZ23]—asserting that the product $\mathbf{A}\mathbf{B}$ of two $n\times n$ real matrices can be computed in $n^{\omega}\sim O(n^{2.37})\ll n^{3}$ time [Str86, Wil12, Le ̵14, DWZ23]—had a profound impact on theoretical computer science and algorithm design. Alas, FMM algorithms are unlikely to be practical on any imaginable hardware, as they inherently rely on recursive divide-and-conquer schemes, which create memory and IO-bottlenecks [KAA20] and lead to highly asymptotic runtimes, earning them the infamous name “galactic algorithms" [LR13]. Another drawback of FMM algorithms in the context of data-driven applications, is that they are agnostic to the input matrices – recursion hinders the algorithm’s ability to exploit structure in the input matrices (e.g., sparsity or invariances), except in a black-box fashion [YZ05].
In attempt to capture practical algorithms, a concurrent line of work has focused on “combinatorial" MatMul algorithms [Arl+70, Wil07, Lin09, Cha15, Yu18, Abb+23], which avoid Strassen-like divide-and-conquer and algebraic decompositions, and have non-asymptotic runtimes [Bal+15, KS17]. While the term “combinatorial" is open to interpretation, it is ultimately meant to capture practical algorithms (more on this below). An outstanding open question in the field is whether combinatorial algorithms can achieve truly subcubic runtime, even for multiplying $n\times n$ boolean matrices :
Conjecture 1.1 ([Abb+23]).
There is no $n^{3-\Omega(1)}$-time combinatorial algorithm for matrix multiplication, even for Boolean matrices.
A recent breakthrough result by Abboud et al. [Abb+23] presented a combinatorial Boolean MatMul algorithm running in $O(n^{3}/2^{\Omega(\sqrt[7]{\log n})})$ time, highlighting the subtlety of the conjecture. However, for real matrices, the best known speedup using combinatorial techniques is much weaker, $O(n^{3}/\log^{2}n)$ [Wil07].
Bypassing 1.1 in pursuit of more practical algorithms, has initiated a long line of research on approximate matrix multiplication (AMM), which studies the best speed-accuracy tradeoff achievable by “combinatorial" algorithms in close to quadratic time. More precisely, for a prescribed parameter $r<n$, the goal is to produce, in $O(rn^{2})$ time, a matrix $\mathbf{C}\in\mathbb{R}^{n\times n}$, which $\varepsilon$-approximates $\mathbf{A}\mathbf{B}$ in the Frobenius norm, where the (normalized) error $\varepsilon\rightarrow 0$ as $r\rightarrow n$. Essentially all known AMM algorithms use randomized sketching or sampling techniques [Sar06, DKM06, Pag13, MZ10, CW13, CNW16, CL99], and the state-of-art after more than 20 years of research is a linear speed-accuracy tradeoff:
$\|\mathbf{C}-\mathbf{A}\mathbf{B}\|^{2}_{F}\leq O\left(\min\left\{\frac{1}{r}\left\lVert\mathbf{A}\right\rVert_{F}^{2}\cdot\left\lVert\mathbf{B}\right\rVert_{F}^{2}\;,\;\frac{n}{r}\|\mathbf{A}\mathbf{B}\|^{2}_{F}\right\}\right),$ (AMM error)
where $\mathbf{C}$ must be produced in $O(rn^{2})$ (randomized) time. We usually refer to the normalized error, given by $\frac{\left\lVert\mathbf{C}-\mathbf{A}\mathbf{B}\right\rVert_{F}^{2}}{\left\lVert\mathbf{A}\right\rVert_{F}^{2}\cdot\left\lVert\mathbf{B}\right\rVert_{F}^{2}}$, where the above bound becomes $1/r$ (we call this the “linear bound”). The first error term can be obtained via the standard “sketch-and-solve" algorithm (e.g., CountSketch [CW13]), while the second one is obtained by a clever output-sensitive variation of CountSketch, using FFT (TensorSketch [Pag13]). The two bounds in are generally incomparable, but they coincide for the (hardest known) distribution of random Gaussian matrices. A conceptual limitation of all aforementioned AMM algorithms (except [CL99] which applies only to nonnegative matrices), is that they use compression techniques, i.e., produce a low-rank or sparse output matrix $\mathbf{C}$. In data-driven applications, most notably LLM training and inference, this is a severe limitation since compression crucially decreases the number of trainable parameters (see [Ail+25] and references therein). Moreover, a recent result of [OP24] proves that for compression-based algorithms, in a setup where Alice and Bob can send $O(rn)$ bits representing their respective input matrices $\mathbf{A},\mathbf{B}\in\mathbb{R}^{n\times n}$ to a “referee" who must then compute the output based on their messages, the error bound in (AMM error) is tight for random Gaussian (or $\pm 1$) matrices.
Our work challenges 1.1, and sheds further light on the power of convolutions (which are omnipresent in practical algorithms) in the context of matrix-multiplication. We develop a new algorithmic framework for approximate matrix-multiplication, which relies only on several invocations of vanilla FFTs after small preprocessing, leading to improvements on the state-of-art in both exact and approximate regimes.
1.1 Our Results
Our work can be viewed as a variation of the group-theoretic framework of Cohn and Umans [CU03]— originally developped for fast matrix multiplication—to the approximate and combinatorial setting, i.e., restricted to abelian groups (polynomials), and disallowing recursive calls [^1] One could argue that FFT is itself a recursive algorithm, however, the divide-and-conquer used in FFT differs fundamentally from the recursion used in FMM algorithms – FFT is used only for the marginal encoding of $\mathbf{A}$ and $\mathbf{B}$ respectively, and not for the multiplication itself (as in FMM), see [Blä13]. Moreover, FFT is clearly used in practice, which is what “combinatorial” algorithms ultimately attempt to capture.: All the algorithms we present rely on fast polynomial multiplication, where our key departure from the Cohn-Umans setting is applying convolutions on low-degree approximations of the polynomials proposed by [Coh+05], while avoiding the use of the asymptotic sum inequality [Sch81]. While our work is merely an initial step toward practical MatMul  algorithms, we provide empirical evidence that our algorithm improves substantially on the best known speed-accuracy tradeoff of approximate matrix multiplication, making it suitable for data-driven applications such as inference speedup in deep neural networks [Ail+25].
Our first result is an exact “combinatorial" MatMul  algorithm, that runs in truly subcubic time:
Theorem 1.2 (Faster Combinatorial MM, Informal).
There is an exact, deterministic algorithm for computing the product of two $n\times n$ matrices $\mathbf{A}\mathbf{B}$, in $O(n^{2.89})$ time. The algorithm only aggregates several convolutions in $\mathbb{Z}_{m}^{k}$ which are computed via FFT.
The algorithm bootstraps a certain construction first proposed by [Coh+05] as a base tensor for FMM. In principle, the algorithm refutes 1.1, but it is unforuntately still highly asymptotic, beating naïve MatMul  ($\sim n^{3}$) only for very large matrices ($n\gtrsim 18^{12}$). This is our main motivation for focusing on the approximate convolutions, where the hope is to exploit the structure of the polynomials used in Theorem 1.2, in order to reduce the complexity of computing their convolution, possibly with some small accuracy loss.
As a proof-of-concept, we first show that a black-box combination of the exact combinatorial algorithm from Theorem 1.2 together with linear sketching, already yields the first super-linear speed-accuracy tradeoff for AMM, improving the state-of-art bound (AMM error):
Theorem 1.3 (Faster AMM via Sketching Convolutions, Informal version of Theorem 7.2).
For any $r\leq n$, there exists a randomized algorithm that, given two square $n\times n$ matrices $\mathbf{A},\mathbf{B}$, produces in time $\tilde{O}(n^{2}r)$ a matrix $\mathbf{C}$, such that w.h.p, $\|\mathbf{C}-\mathbf{A}\mathbf{B}\|_{F}^{2}=O(\frac{1}{r^{1+\delta}}\left\lVert\mathbf{A}\right\rVert_{F}^{2}\left\lVert\mathbf{B}\right\rVert_{F}^{2})$ where $\delta\geq 0.1$.
We reiterate that FFT was already used in the AMM line of work (TensorSketch [Pag13]), albeit in a different way. Our work is the first to surpass the linear $1/r$ speed-accuracy tradeoff (AMM error) for AMM, without using FMM.
One suboptimal feature of Theorem 1.3 is that it is agnostic to the structure of the polynomials being convolved (the polynomials used in Theorem 1.2). The technical centerpiece of our work is a Fourier-concentration lemma for the CKSU polynomials, which leads to a non-blackbox AMM algorithm with substantially better tradeoff, for a large class of product distributions:
Theorem 1.4 (Fourier Truncation Algorithm, Informal version of Theorem 9.6, 9.13).
Let $\mathbf{A},\mathbf{B}$ be two independent $n\times n$ matrices whose entries are i.i.d. from sub-Gaussian distributions $\mathcal{D}_{A},\mathcal{D}_{B}$ with non-zero variance and second moments (e.g., Bernoulli, Rademacher, Gaussian). Then for any $r\leq n/2$, there exists a deterministic algorithm running in time $O(rn^{2})$ that computes a convolution of $3$-variate polynomials, outputting a matrix $\mathbf{C}$, such that $\mathbb{E}[\frac{\left\lVert\mathbf{C}-\mathbf{A}\mathbf{B}\right\rVert_{F}^{2}}{\left\lVert\mathbf{A}\right\rVert_{F}^{2}\cdot\left\lVert\mathbf{B}\right\rVert_{F}^{2}}]\leq O\left(\frac{(n-2r)^{3}}{n^{4}}\right)$. In particular, for random $n\times n$ Gaussian or Rademacher matrices, the expected squared error is
$O\left(\left(\frac{n-2r}{n}\right)^{3}\right)\cdot\left\lVert\mathbf{A}\mathbf{B}\right\rVert_{F}^{2}.$
The constants hidden in the $O(\cdot)$ notation depend on the distributions $\mathcal{D}_{A},\mathcal{D}_{B}$.
For calibration, notice that outputting the trivial expected solution $\mathbf{C}=\mathbb{E}[\mathbf{A}\mathbf{B}]=0^{n\times n}$ without even reading the inputs, already attains a squared error of $\left\lVert\mathbf{A}\mathbf{B}\right\rVert_{F}^{2}$, and this is consistent with Theorem 1.4 when setting $r=0$. For Gaussian matrices, the error bound of Theorem 1.4 is smaller (by a factor of $\frac{(n-r)^{2}}{n^{2}}$) than the best rank-$r$ approximation of the output matrix $\mathbf{A}\mathbf{B}$, obtained by $r$-$\mathsf{SVD}(\mathbf{A}\mathbf{B})$ [MP67] (see Figure Figure 5 for empirical validation). Note that even the fastest algorithms for approximating $r$-$\mathsf{SVD}(\mathbf{A}\mathbf{B})$ to $\varepsilon$-additive accuracy require $\tilde{O}(rn^{2}/\sqrt{\varepsilon})$ time [MM15], via iterated matrix-vector (or block) multiplications ($\mathbf{A}\mathbf{B}x$), which makes these algorithms rather impractical. By contrast, the algorithm in Theorem 1.4 is independent of $\varepsilon$ or the spectral gap $(\sigma_{r+1}(\mathbf{A}\mathbf{B})/\sigma_{r}(\mathbf{A}\mathbf{B}))$, and is parallel (has depth $O(\log n)$ of FFT as opposed to depth $\tilde{O}(r/\sqrt{\varepsilon})$ of iterative Krylov-subspace algorithms). We present empirical evidence (see Figure 5) that corroborates this. We are not aware of non-iterative methods that can achieve $r$-$\mathsf{SVD}(\mathbf{A}\mathbf{B})$ even for random i.i.d matrices.
While the algorithm in Theorem 1.4 cannot be generalized to worst-case matrices (we give a simple counter-example, for which this method cannot work), finding the most general form of this method and adding randomness still remains a fascinating question with potential implications in practice.
A Lower Bound for Direct Convolution Algorithms
We complement our results by proving the limitations of using a single (exact) convolution for AMM. We prove that a single convolution over small abelian groups (e.g. low-degree polynomial multiplication) cannot surpass the linear speed-accuracy tradeoff in (AMM error). We also provide a simple algorithm which matches this error bound.
Theorem 1.5 (AMM Convolution Lower Bound, Informal version of Theorem 8.6, Theorem 8.7).
Any (randomized) algorithm computing the product of two Rademacher matrices $\mathbf{A},\mathbf{B}\in_{R}\left\{\pm 1\right\}^{n\times n}$ using a single convolution $\mathbf{C}_{G}(\mathbf{A},\mathbf{B})$ in an abelian group $|G|\leq rn^{2}$, has expected error $\left\lVert\mathbf{C}_{G}(\mathbf{A},\mathbf{B})-\mathbf{A}\mathbf{B}\right\rVert_{F}^{2}\geq\Omega(\frac{1}{r}\left\lVert\mathbf{A}\right\rVert_{F}^{2}\left\lVert\mathbf{B}\right\rVert_{F}^{2})$. We provide a simple algorithm (degree-$rn^{2}$ univariate polynomials) matching this bound.
Theorem 1.5 generalizes the $\Omega(n^{3})$ lower bound of [CU03] on the size of abelian groups realizing exact matrix multiplication using convolutions, and provides an alternative, elementary algorithm achieving (AMM error), while possibly having better theoretical qualities such as high output rank.
2 Technical Overview
Below we provide a high-level streamlined overview of the ideas behind the proofs of the main results.
Matrix Multiplication via Convolutions
The idea of using group convolutions for computing Matrix Multiplication was first presented in the seminal work of Cohn and Umans [CU03]. Instead of directly multiplying $\mathbf{A},\mathbf{B}\in\mathbb{R}^{n\times n}$, they proposed embedding the matrices as polynomials in $\mathbb{R}[x]_{/(x^{n^{3}}-1)}$ (identify $x^{n^{3}}$ with $1$), choosing a specific set of monomials. In particular, set $P_{\mathbf{A}}(x)=\sum_{i,j\in[n]}\mathbf{A}_{i,j}x^{-n^{2}i+nj}$ and $P_{\mathbf{B}}(x)=\sum_{j,k\in[n]}\mathbf{B}_{j,k}x^{k-nj}$ (all powers are taken modulo $n^{3}$). Let $P_{\mathbf{C}}=P_{\mathbf{A}}\cdot P_{\mathbf{B}}$ denote the polynomial product. Suppose $P_{\mathbf{C}}(x)=\sum_{s\in[n^{3}]}c_{s}x^{s}$, then it is easy to see that
$c_{k-n^{2}i}=\sum_{j\in[n]}\mathbf{A}_{i,j}\mathbf{B}_{j,k}=(\mathbf{A}\mathbf{B})_{i,k}.$ (2.1)
Hence, by computing a single polynomial multiplication via the Fast Fourier Transform (FFT), which is equivalent to the cyclic convolution of the polynomials’ coefficient vectors, one can recover the product matrix $\mathbf{A}\mathbf{B}$. This is true because the group-embedding used in (2.1) satisfies the property that, for any $(i,k)$, the equation $(k^{\prime}-k)+n(j-j^{\prime})+n^{2}(i-i^{\prime})=0$ has exactly $n$ solutions, given by $i^{\prime}=i,k^{\prime}=k$ and $j^{\prime}=j\in[n]$. In other words, only the desired “diagonal" terms $\mathbf{A}_{ij}\mathbf{B}_{jk}$ in the convolution are aggregated to the monomial $x^{k-n^{2}i}$, whereas the undesired “cross-terms" ($i^{\prime}\not=i,k^{\prime}\not=k,j\neq j^{\prime}$) are aggregated into other “garbage" monomials which we simply ignore when recovering $\mathbf{C}$ from $c$. Cohn and Umans termed this property the Triple Product Property (TPP).
Unfortunately, the convolution algorithm in (2.1) is quite useless, as the polynomial $P_{\mathbf{A}}$ has degree $n^{3}$, hence the convolution (via FFT) requires $\Omega(n^{3}\log n)$ time, slower than naïve matrix multiplication(!). As such, [CU03] generalized the group-convolution approach (2.1) and the TPP property to general groups, where instead of polynomials, the matrices $\mathbf{A}$ and $\mathbf{B}$ are embedded into the “group algebra" $\mathbb{C}[G]$ of some general but fixed group $G$, and the two embeddings are then convolved inside $\mathbb{C}[G]$ to produce an element of $\mathbb{C}[G]$ containing the desired output $\mathbf{A}\mathbf{B}$. [^2] More formally, embed the indices $\mathbf{A}(i,j),\mathbf{B}(j^{\prime},k)\to G$ and compute $(\sum_{g\in G}\mathbf{A}_{g}g)\cdot(\sum_{h\in G}\mathbf{B}_{h}h)=\sum_{f\in G}(\sum_{gh=f}\mathbf{A}_{g}\mathbf{B}_{h})f$. See section 3 and [CU03] for more details. The TPP condition for general groups then becomes:
Definition 2.1 (TPP).
We say that a finite group $G$ satisfies the TPP for $\left\langle{n,m,p}\right\rangle$ if there exist sets $S,T,U\subseteq G$ such that $\left|S\right|=n,\left|T\right|=m,\left|U\right|=p$ and for every $s_{1},s_{2}\in S,t_{1},t_{2}\in T,u_{1},u_{2}\in U$ the equality $(s_{1}s_{2}^{-1})(t_{1}t_{2}^{-1})(u_{1}u_{2}^{-1})=1_{G}$ implies $s_{1}=s_{2},t_{1}=t_{2},u_{1}=u_{2}$.
The sets $S,T,U$ play the role of indexing sets: $S$ is used to index the rows of $\mathbf{A}$ and $\mathbf{A}\mathbf{B}$, $T$ the columns of $\mathbf{A}$ and rows of $\mathbf{B}$, and $U$ the columns of $\mathbf{B}$ and $\mathbf{A}\mathbf{B}$. Another view of this definition, is that it captures when the tensor of the group algebra of $G$ (which describes the convolution operator in $\mathbb{C}[G]$) can be restricted to the matrix multiplication tensor [Blä13].
The aforementioned $\Omega(n^{3})$ lower bound on the degree of TPP polynomials in (2.1) is not a coincidence – A simple proof [CU03] shows that any abelian group satisfying the TPP must have $\left|G\right|\geq n^{3}$, which seems to make abelian groups useless. In fact, our Theorem 1.5 generalizes this lower bound to the approximate setting, asserting that convolutions over abelian groups of size $|G|\leq rn^{2}$ cannot yield better than linear (normalized) Frobenius error $\Omega(1/r)$. The proof uses an elegant Cauchy-Schwarz argument on a carefully-chosen set of multiplicities (quadruplets). We show this is tight by providing a very simple and practical algorithm, using a single convolution in $\mathbb{Z}_{rn^{2}}$. The analysis is elementary, in contrast to sketching algorithms [Sar06, DKM06, CW13] which rely on the JL-moment property and Minkowski’s inequality to achieve (AMM error).
The TPP lower bound for abelian groups shifted the focus of the group-theoretic-FMM line of work to finding non-abelian TPP groups of sub-cubic size [^3] In fact, an example of a group $G$ satisfying the TPP with $\left|G\right|=n^{2+o(1)}$ was given in [CU03]. [CU03, Coh+05]. Alas, non-abelian groups generally lack fast convolution algorithms, due to: (i) the absence of a fast Fourier transform algorithm; (ii) The fact that Fourier transform of non-abelian groups transforms convolutions into block-diagonal matrix products, as opposed to element-wise product in the abelian case (see section 3). This is where divide-and-conquer comes into play: applying the algorithm recursively on tiles (a-la Strassen) amortizes the cost of any linear transformation done before or after the recursive step, so applying the Fourier transform of $G$ is asymptotically free of charge. This is also where our approach departs from this line of work – Since we aim for non-recursive (“combinatorial”) algorithms, we must inevitably work with abelian groups which have FFTs. While this seems like a dead-end in light of the $\Omega(n^{3})$ degree lower bound for TPP polynomials, it turns out that multivariate polynomials (e.g., convolutions in $\mathbb{Z}_{m}^{k}$) satisfy a more general TPP condition, which allows to multiply block-matrices faster than the trivial time. We explain this idea, and how we use it to prove Theorem 1.2 next.
2.1 Exact Matrix Multiplication (Proof of Theorem 1.2)
Multivariate Polynomials and Simultaneous TPP
A followup work of Cohn, Kleinberg, Szegedy and Umans [Coh+05], proposed a simple generalization of the (TPP)., termed the Simultaneous Triple Product Property ( (STPP).), which captures the “capacity" of a group $G$ to compute $k$ independent matrix products using a single convolution. (Information-theoretically, this makes sense, since the TPP group-embedding e.g. (2.1) generates extremely sparse polynomials, so intuitively there’s plently of room to “pack" more information into the convolution, see Figure 1). Simply stated, the (STPP). describes when $G$ contains a collection of $d$ triples $(S_{i},T_{i},U_{i})_{i\in[d]}$ (for simplicity, assume $|S_{i}|=|T_{i}|=|U_{i}|=q$ are all equally sized), s.t (i) Each triplet marginally satisfies (TPP)., and (ii) The convolution in $G$ “doesn’t mix" the $d$ triplets ($(t_{j}-s_{j}^{\prime})+(u_{k}-t_{k}^{\prime})=(u_{i}^{\prime}-s_{i})=0$ implies $i=j=k\in[d]$, stated in additive notation). Hence, assuming $G$ has a fast Fourier transform, this property allows to compute $d$ independent $q\times q$ MatMuls for the price of a single convolution $\tilde{O}(|G|)$. Thus, if $|G|\ll dq^{3}$, this is faster than the naïve runtime. Unlike (TPP)., [Coh+05] show that there are nontrivial abelian groups satisfying (STPP)., and provide a simple non-trivial construction over the group $\mathbb{Z}_{m}^{3N}$ ($3N$-variate polynomials with bounded degree $m$ on each variable) with $d=2^{N}$ and $q=(m-1)^{N}$. This basic construction, which we’ll call “CKSU polynomials”, will play a central role in the proof of Theorem 1.4. In its most basic form, by setting $N=1$ (so d=2,q=(m1)d=2,q=(m-1)), it allows embedding two independent matrix products $\mathbf{A}_{1}\times\mathbf{B}_{1}$ and $\mathbf{A}_{2}\times\mathbf{B}_{2}$, of size $m-1\times m-1$ via the convolution of the following polynomials in $\mathbb{Z}_{m}^{3}$:
$\mathbf{A}_{1}=(\mathbf{A}_{i,j}^{(1)})_{i,j=0}^{m-2},\ \mathbf{A}_{2}=(\mathbf{A}_{i,j}^{(2)})_{i,j=0}^{m-2}\quad$ $\leadsto\quad a(x,y,z)=\sum_{i,j=0}^{m-2}\left(\mathbf{A}_{i,j}^{(1)}\cdot x^{i+1}y^{j+1}+\mathbf{A}_{i,j}^{(2)}\cdot y^{i+1}z^{j+1}\right),$ (2.2)
$\mathbf{B}_{1}=(\mathbf{B}_{i,j}^{(1)})_{i,j=0}^{m-2},\ \mathbf{B}_{2}=(\mathbf{B}_{i,j}^{(2)})_{i,j=0}^{m-2}\quad$ $\leadsto\quad b(x,y,z)=\sum_{i,j=0}^{m-2}\left(\mathbf{B}_{i,j}^{(1)}\cdot y^{i+1}z^{j+1}+\mathbf{B}_{i,j}^{(2)}\cdot x^{i+1}z^{j+1}\right).$ (2.3)
Since [Coh+05] aim for the best-possible $n^{\omega}$, they invoke the STPP construction (2.2) recursively using the Asymptotic Sum Inequality [Sch81]. Our simple observation is that one can instead compute the polynomial multiplication (2.2) directly using the group’s FFT (the group is $\mathbb{Z}_{m}^{3N}$) to compute $d$ products of $q\times q$ matrices, asymptotically faster than the naïve algorithm $O(d\cdot q^{3})$ (albeit far from the optimal bound on $\omega$ achievable via recursion). This is illustrated in Figure 1.
Given this black-box algorithm for computing $d$ products of $q\times q$ matrices, we can use it to multiply $dq\times dq$ square matrices $\mathbf{A},\mathbf{B}$ using the following idea. Partition the matrices into $d\times d$ blocks of size $q\times q$. Then computing $\mathbf{A}\mathbf{B}$ requires exactly $d^{3}$ block multiplications (the same as in the scalar case, just with blocks), hence it suffices to invoke the black box algorithm $d^{2}$ times (which can be done in parallel, unlike [Coh+05]). The runtime savings (if any) of a single invocation of the black-box algorithm are thus multiplied by a factor of $d^{2}$. This greatly improves the asymptotics, and substitutes the role of recursion in the analysis. The exact analysis of this idea for the aforementioned construction leads to the result in Theorem 1.2.
Figure 1: The convolution algorithm underlying Theorem 1.2. The group is $\mathbb{Z}_{m}^{3N}$. We are embedding $d=2^{N}$ square matrices of size $q=(m-1)^{N}$ in two polynomials $a,b$, computing their Fourier transforms, and computing the element-wise product. The decoding part is not presented in this figure, but is essentially the same. Invoking this algorithm $d^{2}$ times (possibly in parallel) leads to the fast combinatorial algorithm.
2.2 Approximate Matrix Multiplication via Convolutions
Our main contribution is extending the Cohn Umans framework to the approximate case, i.e., studying approximate analogues of (TPP). and (STPP). sets in abelian groups. Our main motivation here is to mitigate the asymptotics of Algorithm 1 (depicted in Figure 1) and to take another step toward practical MatMul  algorithms, in settings where inexact computation is tolerable (such as Deep Learning [Ail+25], Databases [LLM23], and statistics [Sar06, CW13]). The idea here is to “compress" the CKSU polynomials (2.2) in a way that allows for faster convolution algorithm, without hurting the combinatorial STPP property too much. We provide three low-degree approximation schemes of the (TPP or STPP) polynomials, leading to Theorem 1.3, Theorem 1.4 and Theorem 1.5.
AMM via Sketching Convolutions (Proof of Theorem 1.3)
As a warmup, we show that combining the exact algorithm (1) from Theorem 1.2 with standard linear Sketching ([Sar06, DKM06, CW13]) of the input matrices in a black-box manner, already improves the best known asymptotic speed-accuracy tradeoff (AMM error) achievable by combinatorial algorithms. The proof is straightforward: We first use a random Gaussian matrix $S\in_{R}\mathcal{N}(0,1/\sqrt{n})^{n\times r}$ to “compress" the number of columns (resp. rows) of the input matrices $\mathbf{A}^{\prime}:=\mathbf{A}S$, $\mathbf{B}^{\prime}:=S^{\top}\mathbf{B}$, yielding two $(n\times r)$ matrices. The JL moment property [CW13] then guarantees that
$\mathbb{E}_{S}\|\mathbf{A}\mathbf{B}-\mathbf{A}^{\prime}\mathbf{B}^{\prime}\|_{F}^{2}\leq\frac{1}{r}\left\lVert\mathbf{A}\right\rVert_{F}^{2}\left\lVert\mathbf{B}\right\rVert_{F}^{2}.$ (2.4)
It therefore remains to multiply $\mathbf{A}^{\prime}\mathbf{B}^{\prime}$, which using naïve MatMul  requires $O(rn^{2})$ time, recovering the standard linear speed-accuracy AMM error. Our obvious approach is to use Theorem 1.2 (or more precisely, a rectangular version of it, similar to the convolution in Figure 1), to multiply $\mathbf{A}^{\prime}\mathbf{B}^{\prime}$ slightly faster, in $\approx r^{0.9}n^{2}$ time. Note that the composed algorithm reduces the dimensions, and therefore the degree, of the STPP polynomials being convolved. Since this yields a (small) polynomial saving in runtime, we can afford to increase the number of rows in the sketching matrix $S$ to $\approx r^{1.1}$, while still matching the original runtime budget $O(rn^{2})$. By (2.4) this yields an error of $\approx\frac{1}{r^{1.1}}\left\lVert\mathbf{A}\right\rVert_{F}^{2}\left\lVert\mathbf{B}\right\rVert_{F}^{2}$.
We remark that Pagh’s TensorSketch algorithm [Pag13] also uses FFT (albeit in a different way, to efficiently recover the “heavy-hitters" of an outer product $uv^{\top}\in\mathbb{R}^{n\times n}$), but nevertheless his algorithm still yields only a linear dependence on $r$, see (AMM error).
Sparse Fourier Approximation (Proof of Theorem 1.4)
A suboptimal feature of Theorem 1.3 is that it is agnostic to the structure of the CKSU polynomials (2.2) being convolved, and as an artifact, the sketching algorithm still suffers from the asymptotic caveats of the exact algorithm in Theorem 1.2. The main result of this paper (and also its technical centerpiece) is a non-blackbox low-degree approximation scheme for the CKSU polynomials, based on a surprising Fourier-concentration property of these polynomials. This approximation scheme leads to dramatically better AMM tradeoffs for a large class of product distributions ($\mathbf{A}\perp\mathbf{B}$), both in theory and practice (see section 10). Below we explain the main ideas of this low-degree approximation scheme.
Since convolutions are equivalent to elementwise multiplication in the Fourier domain, Parseval’s identity implies that small Fourier coefficients (frequencies) could be dropped without incurring large (squared) error. Moreover, if most of the energy ($\ell_{2}^{2}$ mass) of the Fourier spectrum of $a,b$ happens to lie on low degree frequencies [^4] In the sense that degree of a character is just the sum of the frequency elements $\xi\in G$ defining it, see section 3., then we obtain a (Fourier basis) low-degree approximation for the polynomials $a,b$, whose convolution is not far (in the $\ell_{2}$ sense) from the exact convolution of the CKSU polynomials $a:=P_{\mathbf{A}},b:=P_{\mathbf{B}}$ in (2.2), which recovers $\mathbf{A}\mathbf{B}$. Truncating both polynomials to the same set of $r$ frequencies (Fourier coefficients), resulting in two new polynomials $a^{\prime},b^{\prime}$, means that we can compute $a^{\prime}*b^{\prime}$ using only $r$ multiplications in the Fourier domain. Thus if we can compute the transform on these frequencies fast, we obtain the desired $\tilde{O}(rn^{2})$ time algorithm in Theorem 1.4. Here, the structure of the subset of retained frequencies comes into play.
For general matrices $\mathbf{A},\mathbf{B}$, expecting the polynomials $\widehat{P_{\mathbf{A}}},\widehat{P_{\mathbf{B}}}$ in (2.2) to be concentrated in the Fourier domain, is too good to be true [^5] For example, a polynomial which has exactly one non-zero monomial will have a perfectly uniform spectrum (uniform in magnitude).. Nevertheless, we prove that for random matrices coming from “nice enough" product distributions $\mathbf{A},\mathbf{B}\sim\mathcal{D}_{A}\times\mathcal{D}_{B}$ (where the marginal entries of $\mathcal{D}_{A},\mathcal{D}_{B}$ are i.i.d), the polynomials $\widehat{P_{\mathbf{A}}},\widehat{P_{\mathbf{B}}}$ are highly concentrated, in the sense that most of the spectrum lies on low-degree coefficients. For the polynomials (2.2), these frequencies turn out to have a very nice structure, namely, they are composed of $3$ “slabs" of the group $\mathbb{Z}_{m}^{3}$ , which can be visualized as a lattice, as in Figure 2 below.
Figure 2: An illustration of the Fourier spectrum of $\mathbb{Z}_{m}^{3}$ viewed as a lattice with $m$ points on each axis. The blue region highlights a single slab of width $r$, as used in Theorem 1.4.
As apparent in (2.2), the polynomial $a:=P_{\mathbf{A}}$ is supported on the monomials $X_{1}=\left\{x^{i}y^{j}\mid 0<i,j\leq m-1\right\}$ and $X_{2}=\left\{y^{i}z^{j}\mid 0<i,j\leq m-1\right\}$. The indicator of the support, $\mathbf{1}_{\mathrm{supp}(a)}$, can be written as $\mathbf{1}_{X_{1}}+\mathbf{1}_{X_{2}}$ (because $X_{1}$ and $X_{2}$ are disjoint). These are polynomials obtained by summing all monomials in $X_{1}$ and $X_{2}$ respectively. It is easily shown that the Fourier transform of $\mathbf{1}_{X_{1}}$ at a frequency $\xi\in\mathbb{Z}_{m}^{3}$ (write $\xi=(\xi_{1},\xi_{2},\xi_{3})$ with $\xi_{i}\in[m]$) is just
$\hat{\mathbf{1}}_{X_{1}}(\xi)=m^{-3/2}\times\begin{cases}(m-1)^{2}&\xi_{1}=\xi_{2}=0,\\ -(m-1)&\xi_{1}=0\text{ xor }\xi_{2}=0,\\ 1&\xi_{1},\xi_{2}\not=0.\end{cases}$
The proof of this fact uses the geometry of complex roots of unity, which are the basic building block of Fourier analysis of finite groups (see section 3). A symmetric result holds for $X_{2}$ and for the indicator of $\mathrm{supp}(b)$. Thus almost all of the Fourier mass of $\mathbf{1}_{\mathrm{supp}(a)}$ resides on facets of the cube, where $\xi_{i}=0$ for some coordinate $i$. The choice of the slabs aims to cover these “heavy” facets. This explains the “diminishing returns” dependence of the error bound in Theorem 1.4 on $r$, as most of the energy is already captured by the thinnest slabs (of width $1$). It is also important to note that this specific structure of slabs lends itself to fast computation of the partial transform and its inverse.
To formalize this intuition, we need to bound the expected ratio between the $\ell^{2}$-mass (energy) outside of the retained frequencies, and the total energy product $\|a\|_{2}^{2}\|b\|_{2}^{2}$. The first step of the proof is a concentration bound of the denominator $\|a\|_{2}^{2}\|b\|_{2}^{2}$, which reduces the problem to bounding the ratio of expectations instead. The concentration bound is just an application of Hoeffding’s inequality, since $\|a\|_{2}^{2}$ and $\|b\|_{2}^{2}$ are just sums of i.i.d. random variables. We pay for this with an added exponentially small error term.
Given this reduction, the proof proceeds as follows. Our goal is to compute the algorithm’s expected non-normalized error. The first step is showing that $\mathbb{E}[\hat{a}(\xi)\cdot\overline{\hat{a}(\eta)}]=C_{a}(\xi-\eta)$ for some function $C_{a}$ which depends on the distribution $\mathcal{D}_{A}$ (similarly, for $b$), i.e., that the above expectation is essentially a function of the difference between the frequencies. The functions $C_{a},C_{b}$ are directly determined by the structure of the polynomials (i.e., their support sets), and can be estimated using a mean and variance part, where the variance part is sensitive to $\delta$ and usually overwhelms the mean part.
The second step, is to express the expected error as a sum of the form $\sum_{\delta\in\mathbb{Z}_{m}^{3}}S\cdot C_{a}(\delta)\cdot C_{b}(\delta)\cdot T(\delta)$, where $S=(m-r)^{3}$ is a constant, and $T(\delta)=\sum_{x\in\mathrm{Output}}\omega^{x\cdot\delta}$ (summing over the output coordinates, from which we read off the result). From here, we continue with a straightforward case analysis and estimate the sum. The obtained upper bound is then used to conclude the theorem.
This bound simplifies for $\left\{0,1\right\}$ random Bernoulli matrices and folded Gaussian matrices (where the values distribute as the absolute value of a Gaussian). This is sufficient to deduce a very strong result for $\left\{\pm 1\right\}$ Rademacher and Gaussian matrices, by splitting the positive and negative parts of the matrices (which distribute like Bernoulli matrices), and using the distributivity of the convolution. In other words, observe that if $a=a_{1}\pm a_{2}$, then $\left\lVert a*b\right\rVert_{2}^{2}\leq 2\left\lVert a_{1}*b\right\rVert_{2}^{2}+2\left\lVert a_{2}*b\right\rVert_{2}^{2}$ (and the same for $b$). The formal details are presented in section 9.
3 Preliminaries
3.1 Fourier Analysis of Abelian Groups
[…]
The group algebra $\mathbb{C}[G]$, can be endowed with an inner product $\left\langle{f,f^{\prime}}\right\rangle=\sum_{g\in G}f(g)\cdot\overline{f^{\prime}(g)}$. The Artin-Wedderburn theorem states that the Fourier transform $\mathcal{F}_{G}$ of $G$ is an isomorphism of algebras that diagonalizes the algebra $\mathbb{C}[G]$ into a direct sum of matrix algebras $\bigoplus_{i}\mathcal{M}_{d_{i}\times d_{i}}(\mathbb{C})$, where $\left\{d_{i}\right\}$ are called the character degrees of $G$. This provides a bound on the tensor rank of the structure tensor of GG [^6] The structure tensor $\mathcal{T}_{G}$ of a group $G$ is a tensor describing the bilinear convolution operator $\mathbb{C}[G]\times\mathbb{C}[G]\to\mathbb{C}[G]$. Formally, we have variables $\left\{X_{g}\right\}_{g\in G},\left\{Y_{g}\right\}_{g\in G},\left\{Z_{g}\right\}_{g\in G}$ and $\mathcal{T}_{G}(X_{g},Y_{h},Z_{f})=1$ if $gh=f$ and 0 otherwise, exactly matching the definition $\delta_{g}*\delta_{h}=\delta_{gh}$. See Appendix B for more details., $\mathcal{T}_{G}$, which is $\sum_{i}d_{i}^{3}$. In case $G$ is abelian, all degrees are $1$, since all linear representations of $G$ are one-dimensional, and so $\bigoplus_{i}\mathcal{M}_{d_{i}\times d_{i}}(\mathbb{C})\cong\mathbb{C}^{\left|G\right|}$, where the multiplication in the latter algebra is just the element-wise product of vectors. In other words, $\mathcal{F}_{G}$ transforms convolution into element-wise multiplication. This is known as the convolution theorem.
Assuming $G$ is abelian, the dual group $\hat{G}$, which is the set of characters (one dimensional representations, i.e., group morphisms $G\to\mathbb{C}^{\times}$), is non-canonically [^7] It requires us to choose a primitive root of unity of order $\left|G\right|$. Our choice will always be $\exp(\frac{2\pi i}{|G|})$. isomorphic to $G$. Thus $\mathbb{C}[G]\cong\mathbb{C}[\hat{G}$], and so the Fourier transform can be viewed as a change of basis, from the $\delta$-functions $\delta_{g}(h)=\mathbf{1}_{g=h}$, to the basis of characters. To ensure it is unitary, we choose the definition of the Fourier transform to be
$f\mapsto(\hat{f}_{\chi})_{\chi\in\hat{G}}=\frac{1}{\sqrt{\left|G\right|}}\left(\left\langle{f,\chi}\right\rangle\right)_{\chi\in\hat{G}}=\left(\frac{1}{\sqrt{\left|G\right|}}\sum_{g\in G}f(g)\chi(g^{-1})\right)_{\chi\in\hat{G}}$.
Under this normalization, we have Parseval’s theorem $\left\lVert f\right\rVert_{2}=\|\hat{f}\|_{2}$, and the convolution theorem is then stated as $\widehat{f*f^{\prime}}=\sqrt{\left|G\right|}\cdot\hat{f}\odot\hat{f}^{\prime}$. The inverse Fourier transform is given by $\hat{f}\mapsto\frac{1}{\sqrt{G}}\left(\sum_{\chi\in\hat{G}}\hat{f}_{\chi}\cdot\chi(g)\right)_{g\in G}$.
By the structure theorem for abelian groups, $G$ can be written as $G=\mathbb{Z}_{n_{1}}\times\ldots\times\mathbb{Z}_{n_{d}}$ for some integers $n_{i},d$. For $n\in\mathbb{N}$, fix $\omega_{n}$ to be a primitive root of unity of order $n$, then $\hat{\mathbb{Z}}_{n}=\left\{\chi_{m}\mid m\in[n]\right\}$ where $\chi_{m}(x)=\omega_{n}^{a\cdot m}$. Moreover, $\hat{G}$ then $\left\{\otimes_{i=1}^{d}\chi_{m_{i}}\mid m_{i}\in[n_{i}]\right\}$, where $\phi\otimes\psi$ is the function $(x,y)\mapsto\phi(x)\cdot\psi(y)$ given that $\phi,\psi$ are functions into $\mathbb{C}$. We define the degree of a character $\chi\in\hat{G}$, given as $\chi=\otimes_{i=1}^{d}\chi_{m_{i}}$, to be $\sum_{i=1}^{d}m_{i}$. The Fourier transform is then explicitly written as
$\hat{f}(\chi_{(m_{1},\ldots,m_{d})})=\frac{1}{\sqrt{\left|G\right|}}\sum_{g\in G}f(g)\cdot\omega_{n_{1}}^{-g_{1}\cdot m_{1}}\cdots\omega_{n_{d}}^{-g_{d}\cdot m_{d}}.$
The tuple $(m_{1},\ldots,m_{d})\in G$ defining $\chi$ is often called the frequency, and so we can drop the $\chi$ notation.
3.2 Paper Organization
In section 4 we present a few constructions we will need for the rest of the paper. In section 5 we present the fast algorithm and its analysis, thus proving Theorem 1.2. In section 6 we present a formal approximation framework based on abelian group convolutions. In section 7 we present the black box Sketch & Solve algorithm, proving Theorem 1.3. In section 8 we present the small abelian group approximation approach, proving Theorem 1.5. In section 9 we present the Fourier sparsification approach behind Theorem 1.4 and prove the theorem. At last, in section 10 we provide some empiric results from our tests, showing supporting evidence for our approach.
4 Constructions
4.1 Vanilla TPP
Let $n,m,p$ be three given integers. Define $G=\mathbb{Z}_{nmp}$, and define
$S_{n,m,p}=\left\{x\cdot mp\mid x\in[n]\right\}\quad,\quad T_{n,m,p}=\left\{x\cdot p\mid x\in[m]\right\}\quad,\quad U_{n,m,p}=\left\{x\mid x\in[p]\right\}.$
Then $(S_{n,m,p},T_{n,m,p},U_{n,m,p})$ realize the (TPP). for $\left\langle{n,m,p}\right\rangle$ in $G$. Let $a,b$ denote the embedded signals of $\mathbf{A},\mathbf{B}$ respectively, then
$\mathrm{supp}(a)=T-S=\left\{xmp+yp\mid x\in[n],y\in[m]\right\}\quad,\quad\mathrm{supp}(b)=U-T=\left\{y-xp\mid x\in[m],y\in[p]\right\},$
showing that $\mathrm{supp}(a)$ is an arithmetic progression (AP) with difference $p$ starting at 0, while $\mathrm{supp}(b)$ is just the interval $([nmp-(m-1)p,nmp-1]\cup[0,p-1])\cap\mathbb{Z}$ (this is a continuous interval because we work on the cyclic group). When $n=m=p$, we denote $S_{n,m,p}=S_{n}$ and similarly for $T,U$.
4.2 CKSU STPP
This construction was first given in [Coh+05] to answer a long-standing question. Set $H=\mathbb{Z}_{m}^{3}$ and denote each copy of $\mathbb{Z}_{m}$ in $H$ by $H_{1},H_{2},H_{3}$ (meaning $H_{1}=\mathrm{Im}(\phi_{1})$ where $\phi_{1}:\mathbb{Z}_{m}\hookrightarrow H$, $x\mapsto(x,0,0)$, and similarly $H_{2},H_{3}$ are defined). Define
$S_{1}^{0}$ $=H_{1}\setminus\left\{0_{H}\right\}$ $S_{2}^{0}$ $=H_{2}\setminus\left\{0_{H}\right\}$ $S_{3}^{0}$ $=H_{3}\setminus\left\{0_{H}\right\}$
$S_{1}^{1}$ $=H_{2}\setminus\left\{0_{H}\right\}$ $S_{2}^{1}$ $=H_{3}\setminus\left\{0_{H}\right\}$ $S_{3}^{1}$ $=H_{1}\setminus\left\{0_{H}\right\}$
The authors in [Coh+05] have shown that this sets satisfy the (STPP).. It allows to compute $2$ independent products of $(m-1)\times(m-1)$ matrices. Using B.3, we set $G=H^{N}$ for some $N\geq 1$, and by a single convolution in $G$ we compute $2^{N}$ independent products of $(m-1)^{N}\times(m-1)^{N}$ matrices. The STPP sets can be indexed by a binary string $\boldsymbol{\alpha}\in\left\{0,1\right\}^{N}$:
$(i\in\left\{1,2,3\right\})\qquad S_{i}^{\boldsymbol{\alpha}}=S_{i}^{\boldsymbol{\alpha}_{1}}\times\ldots\times S_{i}^{\boldsymbol{\alpha}_{N}}.$
5 Sub-Cubic Matrix Multiplication via Convolutions (Proof of Theorem 1.2)
Given an algorithm that computes $k$ products of $n\times n$ matrices, we can use it to multiply $nk\times nk$ matrices $\mathbf{A},\mathbf{B}$, by viewing them as $k\times k$ matrices where each entry is a block of size $n\times n$. Then we need $k^{3}$ block products, which we sum up in a specific order. Thus, it can be computed through $k^{2}$ invocations of the given algorithm. More precisely, let $\mathbf{A}_{i,j},\mathbf{B}_{i,j}$ denote $(i,j)$-th block (so $i,j\in[k]$), then to compute $(\mathbf{A}\mathbf{B})_{i,j}$ we compute the products $\left\{\mathbf{A}_{i,t}\mathbf{B}_{t,j}\mid t\in[k]\right\}$ and sum up the results.
The STPP construction CKSU STPP gives rise to a fast algorithm that computes $k=2^{N}$ products of $(m-1)^{N}\times(m-1)^{N}$ matrices, while costing a single convolution in the group $\mathbb{Z}_{m}^{3N}=H^{N}=G$. Since this group is abelian, it has an FFT, which costs $O(m^{3N}\log m^{3N})$. For completeness, we give a full description of the algorithm in Appendix A.
Remark 5.1.
Using other STPP constructions, like the CW tensor [CW87], leads to faster asymptotic algorithms using the same scheme described above.
Runtime Analysis
For a square matrix of size $n_{0}\times n_{0}$ where $n_{0}=\left(m-1\right)^{N}\cdot 2^{N}$, we can use the notation of the blocking scheme as follows: define $n=\left(m-1\right)^{N}$ and $k=2^{N}$. Write
$\tau=\frac{3\log(m)}{\log(m-1)}\quad,\quad\eta=\frac{\log(2)}{\log(m-1)}.$
Then $n^{\tau}=m^{3N}$, $n^{\eta}=k$, and so $n_{0}=nk=n^{1+\eta}$. Therefore the runtime of the algorithm is
$O(k^{2}m^{3N}\log m^{3N})=O(n^{\tau+2\eta}\log n)=O\left(n_{0}^{\frac{\tau+2\eta}{1+\eta}}\log n_{0}\right).$
For example, setting $m=8$ we obtain $\frac{\tau+2\eta}{1+\eta}\approx 2.89$, and optimizing the expression over $m$, i.e., optimizing $f(x)=\frac{3\log(x)+2\log 2}{\log(x-1)+\log 2}$, we see that it has a local minima at $x\approx 20.23$ for which $f(x)\approx 2.85$, and $\lim_{x\to\infty}f(x)=3$, so the best possible algorithm from this approach has exponent $\approx 2.85$.
Runtime Asymptotics
Suppose the algorithm’s runtime is $C\cdot m^{3N}\log(m^{3N})$ while the naive implementation runtime is $2\cdot 2^{N}\cdot(m-1)^{3N}-2^{N}\cdot(m-1)^{2N}$ (the number of FLOPs). In the table below we give some estimates for the threshold where our algorithm becomes faster, assuming different values of $C$ and $m$. We note that a meticulous calculation of the exact FLOP count for the algorithm would possibly result in bigger constants, which has great effect on the threshold point. Moreover, at these sizes the impracticality of the algorithm stems from the enormous memory usage ($m^{3N}$ for each element in the convolution).
$m$ $C$ minimal value of NN $\text{total matrix size }(n)$
66 22 $30$ $10^{30}$
$10$ 22 $10$ $18^{10}$
$10$ 55 $12$ $18^{12}$
88 55 $16$ $14^{16}$
6 Framework For Approximate MM via Convolutions
Suppose we are given input matrices $\left\{\mathbf{A}_{i}\right\}_{i=1}^{k}$ and $\left\{\mathbf{B}_{i}\right\}_{i=1}^{k}$, all of size $n\times n$. Let
$\mathbf{a}=\begin{bmatrix}\mathrm{vec}(\mathbf{A}_{1})\\ \vdots\\ \mathrm{vec}(\mathbf{A}_{k})\end{bmatrix}\quad,\quad\mathbf{b}=\begin{bmatrix}\mathrm{vec}(\mathbf{B}_{1})\\ \vdots\\ \mathrm{vec}(\mathbf{B}_{k})\end{bmatrix}\quad,\quad\mathbf{c}=\begin{bmatrix}\mathrm{vec}(\mathbf{A}_{1}\mathbf{B}_{1})\\ \vdots\\ \mathrm{vec}(\mathbf{A}_{k}\mathbf{B}_{k})\end{bmatrix}\quad\in\mathbb{R}^{kn^{2}}.$
Problem 6.1 (Group Theoretic Approximation).
Given $r\geq 1$ and possibly $\mathbf{a},\mathbf{b}$, find an abelian group $G$ of size $s$ and matrices $\Pi_{A},\Pi_{B},\Pi_{C}\in\mathbb{C}^{s\times kn^{2}}$, which minimize
$\left\lVert\Pi_{C}^{\top}\mathcal{F}^{-1}_{G}\left(\mathcal{F}_{G}(\Pi_{A}\mathbf{a})\odot\mathcal{F}_{G}(\Pi_{B}\mathbf{b})\right)-\mathbf{c}\right\rVert_{2}^{2},$
under the constraint that the tensor $\mathcal{T}_{G,\Pi_{A},\Pi_{B},\Pi_{B}}$ represented by the (Normal Form). has tensor rank $\mathsf{Trk}(\mathcal{T}_{G,\Pi_{A},\Pi_{B},\Pi_{C}})\leq r$, and preferably can be computed in time $\tilde{O}(r)$. Here $\mathcal{F}_{G}$ denotes the Fourier transform for $G$.
Note that the specific presentation of $\mathcal{T}_{G,\Pi_{A},\Pi_{B},\Pi_{C}}$ with the matrices above might have a large inner dimension (i.e., larger than $r$), and still might have another representation (an equivalent normal form) with inner dimension at most $r$. In other words, $s$ is only an upper bound on the tensor rank. As we will see in later sections, this distinction is necessary to unify our approaches under this formulation.
Moreover, it is important to note that the tensor rank does not directly relate to the runtime of the bilinear algorithm. As highlighted before, recursion makes linear transformations done before and after the recursive step essentially “free”, in the sense that the running time is dominated by the tensor rank. This was first presented by [Str69]. In our setting, we wish to avoid recursion for the algorithm to be combinatorial. Under this view, one needs some extra conditions to hold in order for the runtime to match the tensor rank. Examples for this are low rank linear transformations, or structured linear transformations (FFT and the like). Even without the tensor rank directly determining the runtime, the question remains interesting on its own, as low tensor-rank approximations of the matrix multiplication tensor are not well understood (these are “low-accuracy” approximations).
Of course this problem cannot be truly solved using the current tools, so we fall back to heuristics. It is also important to consider two different variations of this problem:
  1. 1.
    Oblivious Setup: $\mathbf{a},\mathbf{b}$ are unknown. A data-dependent solution will use the input compute the matrices $\Pi_{A},\Pi_{B},\Pi_{C}$ online, with running time that does not exceed the bilinear operation running time. A data-agnostic solution will pre-compute the matrices, meaning they are fixed for every $\mathbf{a},\mathbf{b}$ (this will be useful if we have some extra assumptions, like some distribution which $\mathbf{a},\mathbf{b}$ are taken from).
  2. 2.
    Deep Learning Setup: $\mathbf{a}$ is known (we can always switch the roles of $\mathbf{a},\mathbf{b}$ by computing the transposition), and we can pre-compute the matrices $\Pi_{A},\Pi_{B},\Pi_{C}$ based on $\mathbf{a}$ in advance. This is relevant in deep learning, since the model’s weights are known.
Example 6.2.
For the “full” budget $r=n$ and $k=1$, we can take $G=\mathbb{Z}_{n^{3}}$ (using the Vanilla TPP construction) and set $\Pi_{A},\Pi_{B},\Pi_{C}\in\mathbb{C}^{\left|G\right|\times n^{2}}$ to be the indexing matrices. More precisely, for $X\in\left\{A,B,C\right\}$, we note that each row of $\Pi_{X}$ corresponds to an element of the group, while each column of $\Pi_{X}$ corresponds to a position in the matrix, and so by setting $1$ iff this position is embedded into the element matching the row, and 0 otherwise, we obtain $\Pi_{A}\mathbf{a}=a$ and $\Pi_{B}\mathbf{b}=b$ (here $a,b:G\to\mathbb{R}$ are the signals, viewed as vectors). We give these indexing matrices special notation, by writing $I_{X}:=\Pi_{X}$ as defined above (we will reference them later).
7 Sketch & Solve Approximation
The first approach we will present for solving the approximation problem is to reduce the runtime complexity by sketching the input and then applying the algorithm presented in section 5. This is a “black-box” approach, combining known techniques to obtain a slight improvement. Instead of applying the algorithm on the input matrices, we first sketch the input matrices to lower rank matrices and then compute their product using the algorithm. Note that while the algorithm was presented for square matrices, it can be easily adapted to rectangular matrices, by changing the partitioning scheme. Although this presents no major or practical advancements, it is worth laying out explicitly.
A simple formulation of sketching results (see [Woo14] and references within) for our needs is the following:
Lemma 7.1 (JL-Sketch).
Set $m=\frac{1}{\varepsilon^{2}}$, draw $S\in\mathbb{R}^{n\times m}$ where each coordinate is a random normal Gaussian scaled by $\frac{1}{\sqrt{m}}$. Then
$\mathbb{E}[\left\lVert\mathbf{A}\mathbf{B}-\mathbf{A}SS^{\top}\mathbf{B}\right\rVert_{F}^{2}]=O(\varepsilon^{2}\cdot\left\lVert\mathbf{A}\right\rVert_{F}^{2}\cdot\left\lVert\mathbf{B}\right\rVert_{F}^{2}).$
In fact, this inequality holds with high probability.
In particular, in order to reach an error of $O(\frac{1}{r}\left\lVert\mathbf{A}\right\rVert_{F}^{2}\left\lVert\mathbf{B}\right\rVert_{F}^{2})$ we need $m=r$. Computing $\mathbf{A}S$ and $S^{\top}B$ takes $rn^{2}$ using the naive implementation, and so does computing their product. Note that this process inherently reduces the rank of the output matrix in order to achieve faster computation.
Theorem 7.2.
Suppose we are given two square matrices $\mathbf{A},\mathbf{B}$ of size $n$ and a parameter $r\leq n$ which is not insignificant (compared to $n$). If $n$ is large enough, there exists a sketching algorithm that approximates the multiplication of $\mathbf{A},\mathbf{B}$, by computing a matrix $\mathbf{C}$ in time $O\left(n^{2}\cdot r^{\delta+o(1)}\right)$ where $\delta\leq 0.9$ and $r$ bounds the rank of the matrix $\mathbf{C}$ while also satisfying
$\underset{}{\mathbb{E}}\left[\left\lVert\mathbf{A}\mathbf{B}-\mathbf{C}\right\rVert_{F}^{2}\right]\leq O\left(\frac{1}{r}\left\lVert\mathbf{A}\right\rVert_{F}^{2}\left\lVert\mathbf{B}\right\rVert_{F}^{2}\right).$
In fact, this inequality holds with high probability.
By saying that $n$ and $r$ are large enough, we mean that the combinatorial STPP algorithm achieves savings for $r\times r$ matrices.
Proof.
Fix $m$ and $N$ such that $r=2^{N}(m-1)^{N}$. Write $p=\lceil n/(m-1)^{N}\rceil$ (so $p\geq 2^{N}$ because $n\geq r$). Viewing $S$ as a $p\times 2^{N}$ block matrix, and $\mathbf{A},\mathbf{B}$ as $p\times p$ block matrices (padding them if $p>n/(m-1)^{N}$), the products $\mathbf{A}S,S^{\top}\mathbf{B}$ and $(\mathbf{A}S)(S^{\top}\mathbf{B})$ can all be computed by $p^{2}\cdot 2^{N}$ block products. In each invocation of the STPP algorithm we compute $2^{N}$ block products, so the overall number of algorithm invocations is $p^{2}$. The runtime of each invocation is $O(3Nm^{3N}\log m)$. Thus our overall runtime is
$O(p^{2}\cdot 3N\cdot m^{3N}\log m)=O\left(\frac{n^{2}\cdot 3N\cdot m^{3N}\log m}{(m-1)^{2N}}\right).$
Denoting $s=\frac{3N\cdot m^{3N}\log m}{(m-1)^{2N}}$ we observe that ($\log$ is base $2$)
$\log_{r}(s)=\frac{\log(3N)}{N(1+\log(m-1))}+\frac{3\log(m)-2\log(m-1)}{1+\log(m-1)}.$
The dominating part is $f(m)=\frac{3\log m-2\log(m-1)}{1+\log(m-1)}$ that approaches $1$ from below as $m\to\infty$ and has a minimum at $m\approx 20$ (for which $f(m)\approx 0.85$). For $m\in[8,480]$, we have $f(m)\leq 0.9$. The second part $g(N)=\frac{\log(3N)}{N(1+\log(m-1))}$ vanishes with $N$. Therefore, we can compute the sketch in time $O(n^{2}r^{\delta+o(1)})$, where $\delta\leq 0.9$ (we can always use small $m$ and take $N$ to be large). The claim the follows from the JL sketch theorem. ∎
This theorem shows that for large enough matrices, we can achieve a better bound on the error, in the same runtime, thus beating the state of the art. In the formulation of 6.1, we can write the Sketch & Solve algorithm as follows:
$\Pi_{A}=I_{A}\vec{S}\quad,\quad\Pi_{B}=I_{B}\vec{S^{\top}}\quad,\quad\Pi_{C}=I_{C},$
where $I_{X}$ are the indexing matrices (see 6.2) and $\vec{S}$ is the natural embedding of the linear sketch $S$, seen as a linear morphism $S:\mathcal{M}_{kn\times kn}(\mathbb{R})\to\mathcal{M}_{kn\times r}(\mathbb{R})$, into the space of linear morphisms $\mathbb{R}^{kn^{2}}\to\mathbb{R}^{knr}$ (and similarly $\vec{S^{\top}}$ is the embedding of $S^{\top}$). Therefore $\Pi_{A},\Pi_{B}$ act as linear sketch “compressions” of the input vector into a lower dimensional group.
Remark 7.3.
The Sketch & Solve scheme is rather unique, in the sense that the dimensions of the $\Pi$ matrices correspond to the underlying tensor rank, which is exactly the group size.
8 Low Degree Embedding
The second approach we will present for solving 6.1 is a non-black box approach. The idea is to embed the matrices into sets which are approximately TPP/STPP sets. For the sake of simplicity of presentation we will focus on TPP, but a similar argument could be made for STPP constructions. Relaxing the TPP condition allows for a smaller group size, and therefore assuming the group has an FFT algorithm, we obtain a fast algorithm to compute an approximate matrix multiplication.
8.1 Approximate Triple Product Quantity
We work with square matrix multiplication, although the case of rectangular matrices is very similar, up to some minor adjustments.
Definition 8.1 (Indexing Triplet).
Let $G$ denote a group with size at least $n^{2}$. We say that three subsets $S,T,U\subseteq G$ are an indexing triplet for $G$ and $n$, if $\left|S\right|=\left|T\right|=\left|U\right|=n$ and $\left|S^{-1}T\right|=\left|T^{-1}U\right|=n^{2}$. We denote the collection of indexing triplets by $\mathcal{I}(G,n)$.
Definition 8.2 (Approximate Triple Product Quantity).
Let $G$ be a group of size at least $n^{2}$, then for any indexing triplet $(S,T,U)\in\mathcal{I}(G,n)$ we define the approximate triple product quantity for $(S,T,U)$, denoted by $\rho_{n,G}(S,T,U)$, as the number of solutions to the equation
$s_{1}s_{2}^{-1}t_{1}t_{2}^{-1}=u_{1}u_{2}^{-1}\qquad,s_{1},s_{2}\in S,t_{1},t_{2}\in T,u_{1},u_{2}\in U.$
We then define
$\rho_{n}(G)=\min_{(S,T,U)\in\mathcal{I}(G,n)}\rho_{n,G}(S,T,U),$
i.e. the minimal number of solutions to the equation, and call this quantity the Approximate Triple Product Quantity of $G$ with respect to $n$.
Note that $\rho_{n}(G)\geq n^{3}$, and equality holds if and only if $G$ satisfies the TPP. This follows from the fact that there are always $n^{3}$ trivial solutions to the equation, obtained by setting $s_{1}=s_{2},t_{1}=t_{2}$ and $u_{1}=u_{2}$. Furthermore, a simple argument shows that for an abelian group $G$ to satisfy the TPP for size $n$, we must have $\left|G\right|\geq n^{3}$ (to see this, note that $\phi:S\times T\times U\to G$ given by $(s,t,u)\mapsto s+t+u$ is an injection, using the TPP and the commutativity of $G$).
Remark 8.3.
The (Approximate Triple Product Quantity). can be related to the notion of Efficacy presented in [AZ23]. In particular, letting $\mathcal{T}_{G,S,T,U}$ denote the tensor obtained by the above approximate embedding scheme and convolution, and using the definition of buckets (C.2) presented in Appendix C, we have
$\mathbf{Eff}(\mathcal{T}_{G,S,T,U})^{2}=\sum_{s\in S,u\in U}\frac{n^{2}}{\left|\mathrm{Bucket}_{G}^{s,u}\left(S,T,U\right)\right|}\quad\text{where}\quad\sum_{s\in S,u\in U}\left|\mathrm{Bucket}_{G}^{s,u}\left(S,T,U\right)\right|=\rho_{n,G}(S,T,U).$
In certain constructions, the ATPQ is balanced, in the sense that buckets have almost constant sizes, which implies the (non-squared) efficacy is $\approx n^{4}/\rho_{n,G}(S,T,U)$. Using the efficacy notion might provide a natural generalization to the ATPQ that easily carries over to the (STPP). case.
Lemma 8.4.
Let $G$ be a finite group, $S,A\subset G$, then
$\left|\left\{(x_{1},x_{2},a_{1},a_{2})\in S^{2}\times A^{2}\mid x_{1}x_{2}^{-1}=a_{1}a_{2}^{-1}\right\}\right|\geq\frac{\left|A\right|^{2}\cdot\left|S\right|^{2}}{\left|G\right|}.$
Proof.
Note that $x_{1}x_{2}^{-1}=a_{1}a_{2}^{-1}$ if and only if $a^{-1}_{1}x_{1}=a_{2}^{-1}x_{2}$, so we can rewrite the LHS by -
$\sum_{g\in G}$ $\left|\left\{(x_{1},x_{2},a_{1},a_{2})\in S^{2}\times A^{2}\mid a_{1}^{-1}x_{1}=g=a_{2}^{-1}x_{2}\right\}\right|$
$=\sum_{g\in G}\left|\left\{(x,a)\in S\times A\mid a^{-1}x=g\right\}\right|^{2}=(*)$
By Cauchy-Schwarz we have
$(*)\geq\frac{1}{\left|G\right|}\left(\sum_{g\in G}\left|\left\{(x,a)\in S\times A\mid a^{-1}x=g\right\}\right|\right)^{2}.$
The proof is complete by noting
$\sum_{g\in G}\left|\left\{(x,a)\in S\times A\mid a^{-1}x=g\right\}\right|=\left|A\right|\cdot\left|S\right|.$
Corollary 8.5.
Let $G$ be a finite abelian group, then the (Approximate Triple Product Quantity). satisfies
$\rho_{n}(G)\geq\frac{n^{6}}{\left|G\right|}.$
Proof.
Let $S,T,U\subset G$ be subsets of size $n$. Recall that $\rho_{n,G}(S,T,U)$ is the number of solutions to the following equation, using additive notation,
$s_{1}-s_{2}+t_{1}-t_{2}+u_{1}-u_{2}=0,$
which is equivalent to (by commutativity),
$s_{1}-s_{2}=(t_{2}-u_{1})-(t_{1}-u_{2}).$
Denote $A=TU^{-1}=T-U=\{t-u\mid t\in T,u\in U\}$, and apply 8.4) on $S,A$ to obtain that the number of solutions to the last equation (equal to $\rho_{n,G}(S,T,U)$) is at least $\left|A\right|^{2}\cdot\left|S\right|^{2}/\left|G\right|$. We finish by noting that $\left|A\right|$ must equal $n^{2}$, or more precisely that $n^{2}=\left|T^{-1}U\right|=\left|TU^{-1}\right|$ by definition of an indexing set, due to the fact that $t_{1}u_{1}^{-1}=t_{2}u_{2}^{-1}$ if and only if $u_{1}t_{1}^{-1}=u_{2}t_{2}^{-1}$ if and only if (using the fact $G$ is abelian) $t_{1}^{-1}u_{1}=t_{2}^{-1}u_{2}$. ∎
For an abelian group with $\left|G\right|=rn^{2}$, we conclude that $\rho_{n}(G)\geq\frac{n^{4}}{r}$. Note that the corollary doesn’t hold for non-abelian groups, because the proof uses commutativity to obtain the expression used in 8.4.
8.2 PolyForm Algorithm
Given an indexing triplet $S,T,U\subset G$, we index $\mathbf{A},\mathbf{B}$ using these sets (while implicitly defining an order on $S,T,U$) and embed the matrices into $\mathbb{C}[G]$, by setting $a_{s^{-1}t}=\mathbf{A}_{s,t}$ and $b_{t^{-1}u}=\mathbf{B}_{t,u}$. We then compute their convolution c=abc=a*b and decode $\mathbf{C}_{s,u}=c_{s^{-1}u}$, which takes $O(\left|G\right|\log\left|G\right|)$ assuming $G$ has an FFT algorithm (for example, if $G$ is abelian). The (Approximate Triple Product Quantity). determines the number of noise terms, which we call “collisions”, since
$c_{s_{1}^{-1}u_{1}}=\sum_{s_{2},t_{1},t_{2},u_{2}:\ s_{2}^{-1}t_{1}t_{2}^{-1}u_{2}=s_{1}^{-1}u_{1}}a_{s^{-1}_{2}t_{1}}b_{t_{2}^{-1}u_{2}},$
and a term is summed into $\mathbf{C}$ if and only if it is a solution to the equation $s_{1}s_{2}^{-1}t_{1}t_{2}^{-1}=u_{1}u_{2}^{-1}$. There are $n^{3}$ valid solutions, which are $s_{2}=s_{1},t_{2}=t_{1},u_{2}=u_{1}$, which encode the matrix multiplication, while all other terms are noise terms. In the formulation of 6.1, the matrices $\Pi_{X}$ are just indexing matrices $I_{X}\in\mathcal{M}_{\left|G\right|\times n^{2}}$ which embed into the sets $S^{-1}T,T^{-1}U,S^{-1}U$, and the rank of the resulting tensor is just the tensor of the structure tensor of $G$.
The actual Frobenius error of this algorithm then depends on the squared sum of the coefficients corresponding to these error terms. Since the collisions are pre-determined by the indexing triplet, they might be skewed, meaning that certain positions ($s^{-1}u$) might suffer from relatively large numbers of collisions. Moreover, big error coefficients might end up together under certain orderings of $S,T,U$, but not under different orderings. To mitigate the first effect, we should be mindful of how skewed a certain indexing triplet is. Moreover, to mitigate the latter effect, we can randomly choose the ordering of the indexing sets.
Another randomness we can add to the algorithm is adding structured random signs to $\mathbf{A},\mathbf{B}$. The goal here is to promote error terms cancellations, as we’ll show next. Take $\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{\gamma}$ to be independent random vectors $\left\{\pm 1\right\}^{n}$ which are drawn uniformly at random. Define $\tilde{\mathbf{A}}=(\boldsymbol{\alpha}\boldsymbol{\gamma}^{\top})\odot\mathbf{A}$ and $\tilde{\mathbf{B}}=(\boldsymbol{\gamma}\boldsymbol{\beta}^{\top})\odot\mathbf{B}$, where $\odot$ denotes the Hadamard product of matrices. Then $\tilde{\mathbf{A}}_{s,t}=\boldsymbol{\alpha}_{s}\boldsymbol{\gamma}_{t}\cdot\mathbf{A}_{s,t}$ and $\tilde{\mathbf{B}}_{t,u}=\boldsymbol{\gamma}_{t}\boldsymbol{\beta}_{u}\mathbf{B}_{t,u}$. Since $\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{\gamma}$ are independent, we have that
$\mathbb{E}[(\tilde{\mathbf{A}}_{s,t}\cdot\tilde{\mathbf{B}}_{t^{\prime},u})^{2}]=\begin{cases}(\mathbf{A}_{s,t}\cdot\mathbf{B}_{t,u})^{2}&t=t^{\prime},\\ 0&\text{else.}\end{cases}$
This is exactly what we are interested in – canceling collisions which can be characterized by the condition $t\not=t^{\prime}$.
This defines a randomized oblivious algorithm, which we denote by $\mathrm{RPF}$ (stands for Randomized PolyForm). Our goal is to understand the worst case error of this algorithm, meaning
$\mathcal{E}=\max_{\mathbf{A},\mathbf{B}\not=0}\left\{\frac{\mathbb{E}\left[\left\lVert\mathrm{RPF}(\mathbf{A},\mathbf{B})-\mathbf{A}\mathbf{B}\right\rVert_{F}^{2}\right]}{\left\lVert\mathbf{A}\right\rVert_{F}^{2}\cdot\left\lVert\mathbf{B}\right\rVert_{F}^{2}}\right\}.$
Note that the algorithm implicitly uses some indexing triplet, thus we define $\mathcal{E}_{G}=\min_{(S,T,U)\in\mathcal{I}(G,n)}\mathcal{E}_{S,T,U}$ as the best algorithm error achieved by the group $G$. The error can be related with the (Approximate Triple Product Quantity)., as described in the next theorem. The proof is given in Appendix C. The general idea of the proof is using Yao’s principle with a distribution over matrices with i.i.d. values in $\left\{\pm 1\right\}$, and showing that in expectation the RPF error is related with $\rho_{n}(G)$.
Theorem 8.6.
If $G$ is a finite group with $\left|G\right|\geq n^{2}$, then $\mathcal{E}_{G}=\Omega\left(\frac{\rho_{n}(G)-n^{3}}{n^{4}}\right)$.
8.3 Tight Results In The Abelian case
Next, we show that for a specific abelian group $\mathbb{Z}_{rn^{2}}$, where $r\leq n$, there is a setup achieving this error. The proof is provided in Appendix C.
Theorem 8.7.
For $G=\mathbb{Z}_{rn^{2}}$, where $r\leq n$, there exist an indexing triplet $(S,T,U)$ based on arithmetic progressions, such that $\mathcal{E}_{(S,T,U)}=O\left(\frac{\rho_{n}(G)-n^{3}}{n^{4}}\right)$. The runtime of PolyForm in this setup in $O(rn^{2}\log n)$.
It is worth laying out the construction explicitly. The indexing triplet is given by
$S=\left\{rn\cdot i\mid i\in[n]\right\},\ T=\left\{(r-1)\cdot i\mid i\in[n]\right\},\ U=\left\{r\cdot i\mid i\in[n]\right\}.$
The set A=(S+U)(S+U)A=(S+U)-(S+U) is the arithmetic progression with difference $r$, and $t-t^{\prime}\in A$ (for $t,t^{\prime}\in T$) if and only if $t-t^{\prime}=kr(r-1)$ for some $k\in\mathbb{Z}$. Therefore $\rho_{n,\mathbb{Z}_{rn^{2}}}(S,T,U)=n^{3}\cdot\lfloor n/r\rfloor$. By Theorem 8.6 and Theorem 8.7, this is tight for abelian groups using the PolyForm framework.
Remark 8.8.
This simple randomized algorithm recovers state of the art results for AMM achieved using linear sketching techniques. Although the error is the same, there might be other qualitative differences. An important difference in the context of deep learning, is the rank of the output matrix. Our empiric trials suggest that the rank of the output matrices in this approach is generally full, even for small $r$. Compared with sketching techniques, where $r$ is also the upper bound on the rank, this is a major difference.
9 Fourier Domain Truncation
The third approach we will present for solving 6.1 is based on the following simple idea: Given $r$, choose (randomly or deterministically) a set $K\subset\hat{G}$ (where $G$ is an abelian group) of $r$ frequencies of the Fourier transform of the convolution c=abc=a*b, compute $c^{\prime}:=\mathcal{F}^{-1}_{G}(\hat{c}|_{K})$, and decode $c^{\prime}$, i.e., return $I_{C}^{\top}c^{\prime}$ where $I_{C}$ is the indexing matrix for $\mathbf{c}$.
In order to show this is a useful approach we must show two things:
  1. 1.
    Bounded Error: We need to bound $\left\lVert I_{C}(c^{\prime})-I_{C}(c)\right\rVert_{2}^{2}$ in terms of $\left\lVert a\right\rVert_{2},\left\lVert b\right\rVert_{2}$ and $r$.
  2. 2.
    Low Complexity: We need to show that $\hat{a}|_{K},\hat{b}|_{K}$ and $c^{\prime}$ can be computed in time $\tilde{O}(r)$.
Remark 9.1.
A trivial bound on the runtime complexity of computing $\hat{a}|_{K},\hat{b}|_{K}$ and $c^{\prime}$ is $O(rkn^{2})$. This follows from observation that the algorithm can be written down as follows: Given inputs $\mathbf{a},\mathbf{b}\in\mathbb{R}^{kn^{2}}$ and a set of $r$ frequencies $K\subset\hat{G}$, computing $\mathbf{c}^{\prime}$ is the same as computing
$f(\mathbf{a},\mathbf{b})=I_{C}^{\top}\left(\mathcal{F}_{G}^{-1}\left(P_{K}([\mathcal{F}_{G}(I_{A}\mathbf{a})]\odot[\mathcal{F}_{G}(I_{B}\mathbf{b})]\right)\right),$
where $P_{K}$ is the idempotent projection onto $K$ (meaning $(P_{K}x)_{\chi}=x_{\chi}\cdot\mathbf{1}_{\chi\in K}$). Note that $P_{K}(x\odot y)=P_{K}(x)\odot P_{K}(y)$, and so by pre-computing some of the matrix products (which are fixed), we obtain smaller matrices and vectors, leading the complexity to $O(rkn^{2})$.
In order to achieve low complexity, we will thus need $K$ to admit some extra structure. We will see examples for this later in the section.
9.1 Basic Properties
Let c=abc=a*b denote the convolved signal and $c^{\prime}=\mathcal{F}^{-1}(\hat{c}|_{K})$ denote the obtained signal. We always assume some underlying abelian group $G$, so we drop the notation of $G$ where it is clear. By the convolution theorem, we know that $\hat{c}_{\chi}=\sqrt{\left|G\right|}\cdot\hat{a}_{\chi}\cdot\hat{b}_{\chi}$, and so it is intuitive to think that large magnitude coefficients in $\hat{c}$ must originate from large magnitude coefficients in $\hat{a},\hat{b}$. We would have liked to say that $\left\lVert c-c^{\prime}\right\rVert_{2}\leq\left\lVert\hat{a}|_{K^{c}}\right\rVert_{2}\cdot\|\hat{b}|_{K^{c}}\|_{2}$, but unfortunately this is not correct. A correct statement would be:
Lemma 9.2.
It holds $\left\lVert c^{\prime}-c\right\rVert_{2}^{2}=\left\lVert\hat{c}|_{K}-\hat{c}\right\rVert_{2}^{2}\leq\left|G\right|\cdot\left\lVert\hat{a}|_{K^{c}}\right\rVert_{2}^{2}\|\hat{b}|_{K^{c}}\|_{2}^{2}$.
Proof.
The first equality follows from Parseval’s identity. The inequality follows from viewing $c-c^{\prime}=\mathcal{F}^{-1}(\hat{c}|_{K^{c}})$ as the convolution of two signals $a^{\prime}=\mathcal{F}^{-1}(\hat{a}|_{K^{c}})$ and $b^{\prime}=\mathcal{F}^{-1}(\hat{b}|_{K^{c}})$ (we can do this using the convolution theorem), and so by Young’s inequality (which says $\left\lVert x*y\right\rVert_{2}\leq\left\lVert x\right\rVert_{1}\cdot\left\lVert y\right\rVert_{2}$) and the $\ell^{1}$-$\ell^{2}$ inequality ($\left\lVert x\right\rVert_{1}\leq\sqrt{n}\cdot\left\lVert x\right\rVert_{2}$ for $x\in\mathbb{C}^{n}$), we obtain
$\left\lVert c-c^{\prime}\right\rVert_{2}^{2}=\left\lVert a^{\prime}*b^{\prime}\right\rVert_{2}^{2}\leq\left\lVert a^{\prime}\right\rVert_{1}^{2}\cdot\left\lVert b^{\prime}\right\rVert_{2}^{2}\leq\left|G\right|\cdot\left\lVert a^{\prime}\right\rVert_{2}^{2}\cdot\left\lVert b^{\prime}\right\rVert_{2}\overset{\text{Parseval}}{=}\left|G\right|\cdot\left\lVert\hat{a}|_{K^{c}}\right\rVert_{2}^{2}\cdot\left\lVert\hat{b}|_{K^{c}}\right\rVert_{2}^{2}.$
The $\ell^{1}$-$\ell^{2}$ is tight when the coordinates of the vector are uniform in magnitude, and loose when the coordinates are highly concentrated. So in some cases the bound might be improved upon.
The actual error in this approach is not given by the vector $c^{\prime}-c$, but rather by the projection of this vector onto the embedding set of the output matrix. Suppose $I_{C}$ denotes the indexing matrix of the output, then we care about $\left\lVert I_{C}^{\top}(c^{\prime}-c)\right\rVert_{2}^{2}$. Note that $I_{C}^{\top}$ is a contraction (because it simply drops coordinates), so the above expression is an upper bound of the error. Essentially, this shows that the error can be bounded by $\ell^{2}$-concentration of $\hat{a},\hat{b}$ on the set $K$.
Remark 9.3.
The convolution c=abc=a*b usually has large support (almost $\left|\mathrm{supp}(a)\right|\cdot\left|\mathrm{supp}(b)\right|$), but in general we cannot assume anything about the ratio $\left\lVert I_{C}^{\top}c\right\rVert_{2}/{\left\lVert c\right\rVert_{2}}$. For example, in the Vanilla TPP construction, we can take $\mathbf{A}$ to be all zeros except for $\mathbf{A}_{1,1}=M>0$ and similarly take $\mathbf{B}$ to be all zeros except for $\mathbf{B}_{2,1}=1$. Then $a$ and $b$ are all zeros except for two coordinates, whose product is not a desired term, leading to c=abc=a*b satisfying $I_{C}^{\top}c=0$ while $\left\lVert c\right\rVert_{2}=M$. Thus the ratio may be 0, or very close to it, in case there is some dominating non-desired term. In such cases, an approximation of $c$ might essentially be an approximation of the non-desired terms. This is the regime of the problem where our algorithm under performs, and is more suitable for output-sensitive algorithms, like that of [Pag13, BN20, Lin09].
The Fourier transforms for the different constructions presented here have very rigid structures:
  1. 1.
    CKSU STPP: Consider the case of $N=1$ (for $N>1$ the idea is the same, just with more notation and dimensions), i.e., the group $\mathbb{Z}_{m}^{3}$. We can write $a=a_{1}+a_{2}$ where $a_{1},a_{2}$ correspond to $\mathbf{A}_{1},\mathbf{A}_{2}$ respectively. Then $a_{1}$ is supported on $\left\{(x,y,0)\mid x,y\not=0\right\}$, which implies
    $(\hat{a}_{1})_{\xi}=C\cdot\sum_{x=1}^{m-1}\sum_{y=1}^{m-1}a_{x,y,0}\omega^{\xi_{1}x+\xi_{2}y},$
    (where $\omega=\exp(-2\pi i/m)$) showing that the spectrum of $\hat{a}_{1}$ is actually $2$-dimensional and fixed on the third axis (independent of $k_{3}$). Similarly, $a_{2}$ is supported on $\left\{(0,x,y)\right\}$ and so the spectrum of $\hat{a}_{2}$ is fixed on the first axis. The same idea is true for $b$, just with different axes.
  2. 2.
    Vanilla TPP: Recall that $a$ is supported on an arithmetic progression, with step size $n$. We have
    $\hat{a}_{\xi}=C\sum_{s=0}^{n-1}\sum_{t=0}^{n-1}a_{sn^{2}+nt}\omega^{\xi(sn^{2}+tn)},$
    (where $\omega=\exp(-2\pi i/n^{3})$) thus if $\xi_{1}\equiv\xi^{\prime}_{1}\pmod{n^{2}}$ we obtain that $\omega^{\xi_{1}(sn^{2}+tn)}=\omega^{\xi_{1}^{\prime}(sn^{2}+tn)}$ for every $s$ and $t$. Thus, $\hat{a}$ is $n^{2}$-periodic. Therefore, to estimate the top-$r$ Fourier coefficients of the convolution (assuming $r\geq n^{2}$), it suffices to estimate the top coefficients of $\hat{b}$, because we have an implicit representation of size $n^{2}$ for $\hat{a}$ which can be computed in time $O(n^{2}\log n)$. Sadly, $\hat{b}$ is not periodic. However, in the Deep Learning context, the weights matrix is fixed, and so using this construction one can treat the weights matrix (or its transpose) as $\mathbf{B}$, and compute the full spectrum of $\hat{b}$ offline. Then, assuming it is concentrated, the heavy set can be pre-computed and compared with $\hat{a}$ in real time.
9.2 Concentration In Product Distributions
This subsection is devoted to proving a concentration results for the special case in which we are given matrices whose entries are i.i.d. from a real product distribution. Throughout this subsection we will work with the CKSU STPP construction.
Notation: Let us start by fixing some notation. We are working in the group $G=\mathbb{Z}_{m}^{3}$. Define
$X_{1}=\left\{(x,y,0)\mid x,y\not=0\right\}\quad,\quad X_{2}=\left\{(0,x,y)\mid x,y\not=0\right\}\quad,\quad X_{3}=\left\{(x,0,y)\mid x,y\not=0\right\}.$
Recall that these sets form the STPP sets in the construction. In particular, given matrices $\mathbf{A}_{1},\mathbf{A}_{2},\mathbf{B}_{1},\mathbf{B}_{2},$ we embed $\mathbf{A}_{1}$ on $X_{1}$, $\mathbf{A}_{2}$ and $\mathbf{B}_{1}$ on $X_{2}$, and $\mathbf{B}_{2}$ on $X_{3}$. Let $f_{i}=\mathbf{1}_{X_{i}}$ denote the indicator function of $X_{i}$. Our $a$ signal, is constructed by two signals $a=a_{1}+a_{2}$, such that $a_{1},a_{2}$ are encoding $\mathbf{A}_{1},\mathbf{A}_{2}$, respectively. Then $a$ will be supported on $X_{1}\cup X_{2}$, so we define $f_{a}:=\mathbf{1}_{X_{1}\cup X_{2}}=f_{1}+f_{2}$ (using the fact the set $X_{i}$ are pairwise disjoint). Similarly, $b=b_{1}+b_{2}$ will be supported on $X_{2}\cup X_{3}$, so we define $f_{b}:=f_{2}+f_{3}$. The output coordinates, which we read off to obtain our output, are $X_{1}\cup X_{3}$.
[…]
Given a real sub-Gaussian distribution $\mathcal{D}$ we denote the distribution’s parameters as:
$\mu=\mathbb{E}_{x\sim\mathcal{D}}[x],\quad\nu=\mathbb{E}_{x\sim\mathcal{D}}[x^{2}],\quad\sigma^{2}=\mathsf{Var}(\mathcal{D})=\nu-\mu^{2},$
and we let $p$ denote the variance proxy of the squared distribution, $x^{2}$ for $x\sim\mathcal{D}$. The variance proxy of a distribution $\mathcal{D}$ is the smallest real number to satisfy $\mathbb{E}_{x\sim\mathcal{D}}[e^{\lambda(x-\mathbb{E}x)}]\leq e^{p^{2}\lambda^{2}/2}$ for any $\lambda\in\mathbb{R}$. In case $\mathcal{D}$ is a bounded in $[a,b]$, $p\leq\frac{(b-a)}{2}$ (Hoeffding’s lemma). We assume all the parameters are bounded.
We will also need the following useful fact:
Fact 9.4.
Let $1<n\in\mathbb{N}$, and let $\zeta$ denote a primitive root of unity of order $n$. Then for every $k\in\mathbb{Z}$, it holds:
$\sum_{j=0}^{n-1}\zeta^{j\cdot k}=\begin{cases}0&k\not\equiv 0\pmod{n},\\ n&\text{else.}\end{cases}$
The most basic lemma we need is the spectrum structure of the indicator functions:
Lemma 9.5.
For $i\in\left\{1,2,3\right\}$ and $\xi\in\mathbb{Z}_{m}^{3}$, we have
$\hat{f}_{i}(\xi)=\frac{1}{m^{3/2}}\times\begin{cases}(m-1)^{2}&\xi_{i}=\xi_{i+1}=0,\\ -(m-1)&\xi_{i}=0\text{ xor }\xi_{i+1}=0,\\ 1&\xi_{i},\xi_{i+1}\not=0.\end{cases}$
Proof.
A direct computation shows:
$\hat{f}_{i}(\xi)$ $=\frac{1}{\sqrt{\left|\mathbb{Z}_{m}^{3}\right|}}\sum_{(x,y,z)\in\mathbb{Z}_{m}^{3}}f_{i}(x,y,z)\cdot\omega^{\xi_{1}x+\xi_{2}y+\xi_{3}z}$
$=\frac{1}{m^{3/2}}\sum_{(x,y,z)\in X_{i}}1\cdot\omega^{\xi_{1}x+\xi_{2}y+\xi_{3}z}$
$=\frac{1}{m^{3/2}}\sum_{0\not=x,y\in[m]}\omega^{\xi_{i}x+\xi_{i+1}y}$
$=\frac{1}{m^{3/2}}\left(\sum_{0\not=x\in[m]}\omega^{\xi_{i}x}\right)\left(\sum_{0\not=y\in[m]}\omega^{\xi_{i+1}y}\right)$
$=\frac{1}{m^{3/2}}\left(-1+\sum_{x\in[m]}\omega^{\xi_{i}x}\right)\left(-1+\sum_{y\in[m]}\omega^{\xi_{i+1}y}\right),$
and so the claim follows from 9.4. ∎
9.2.1 Main Theorem Statement
Now, we can state the main concentration theorem, i.e., the exact version of Theorem 1.4:
Theorem 9.6 (Concentration Theorem).
Let $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$ be two real sub-Gaussian distributions, with non-zero variance and second moments. Suppose $\mathbf{A}_{1},\mathbf{A}_{2},\mathbf{B}_{1},\mathbf{B}_{2}\in\mathbb{R}^{m-1\times m-1}$ take independent values in distributions $\mathcal{D}_{A},\mathcal{D}_{B}$. Let $r\leq m$ be a parameter, and define $K=K_{1}\cup K_{2}\cup K_{3}$ where $K_{i}=\left\{\xi\in\mathbb{Z}_{m}^{3}\mid\xi_{i}<r\right\}$. Consider the convolution of the embedded signals $a=a_{1}+a_{2},b=b_{1}+b_{2}\in\mathbb{C}[\mathbb{Z}_{m}^{3}]$ (using the CKSU STPP construction), denoted by c=abc=a*b. Let $c^{\prime}=\mathcal{F}^{-1}(\widehat{a*b}|_{K})$ denote the approximation of $c$. Then the expected normalized error of the algorithm satisfies
$\mathbb{E}\left(\frac{\|I_{C}^{\top}(c^{\prime}-a*b)\|_{2}^{2}}{\left\lVert a\right\rVert_{2}^{2}\left\lVert b\right\rVert_{2}^{2}}\right)\leq O\left(\frac{(m-r)^{3}}{(m-1)^{4}}\cdot\frac{\sigma_{A}^{2}\sigma_{B}^{2}}{\nu_{A}\nu_{B}}\right)+8(m-1)^{2}\cdot\exp(-\gamma(m-1)^{2}),$
where $\gamma=(6-4\sqrt{2})\cdot\min\left(\frac{\nu_{A}^{2}}{p_{A}^{2}},\frac{\nu_{B}^{2}}{p_{B}^{2}}\right)$.
We call the sets $K_{1},K_{2},K_{3}$, slabs of width $r$, as depicted in Figure 2.
9.2.2 Proving the Main Theorem
We start with a basic lemma, whose proof is deferred to Appendix D.
Lemma 9.7 (Low Correlation Lemma).
In the same setup of Theorem 9.6, let $\xi,\eta\in K^{c}$ and $\delta=\xi-\eta$. Then there exist functions $C_{a},C_{b}:\mathbb{Z}_{m}^{3}\to\mathbb{R}$ such that $\mathbb{E}[\hat{a}(\xi)\cdot\overline{\hat{a}(\eta)}]=C_{a}(\delta)$ and $\mathbb{E}[\hat{b}(\xi)\cdot\overline{\hat{b}(\eta)}]=C_{b}(\delta)$. Moreover, we estimate
$\left|C_{a}(\delta)\right|\leq\frac{4\mu_{A}^{2}}{m^{3}}+\sigma_{A}^{2}\cdot\begin{cases}\frac{2}{m}&\delta\in\left\{(0,0,*),(*,0,0)\right\},\\ \frac{2}{m^{2}}&\delta_{1}=0,\delta_{3}=0,\\ \frac{2}{m^{3}}&\text{otherwise,}\end{cases},\left|C_{b}(\delta)\right|\leq\frac{4\mu_{B}^{2}}{m^{3}}+\sigma_{B}^{2}\cdot\begin{cases}\frac{2}{m}&\delta\in\left\{(*,0,0),(0,*,0)\right\},\\ \frac{2}{m^{2}}&\delta_{1}=0,\delta_{2}=0,\\ \frac{2}{m^{3}}&\text{otherwise.}\end{cases}$
Notation. Let us introduce some convenient notation. Write $\mathcal{E}=\sum_{x\in X_{1}\cup X_{3}}\mathbb{E}[|e(x)|^{2}]$, where $e=c^{\prime}-a*b$ from the theorem’s statement. By the Fourier inversion formula,
$\mathcal{E}=\sum_{x\in X_{1}\cup X_{3}}\mathbb{E}\left[\left|\frac{1}{m^{3/2}}\sum_{\xi\in K^{c}}\hat{e}(\xi)\cdot\omega^{x\cdot\xi}\right|^{2}\right]=\frac{1}{m^{3}}\sum_{x\in X_{1}\cup X_{3}}\sum_{\xi,\eta\in K^{c}}\mathbb{E}[\hat{e}(\xi)\cdot\overline{\hat{e}(\eta)}]\cdot\omega^{x(\xi-\eta)}.$
Denote $\delta=\xi-\eta$. By the Convolution Theorem (formulated as $\widehat{a*b}(\xi)=\sqrt{\left|G\right|}\cdot\hat{a}(\xi)\cdot\hat{b}(\xi)$), we know $\hat{e}$ is supported on $K^{c}$, where $\hat{e}(\xi)=\widehat{a*b}(\xi)$, therefore by 9.7 we conclude that
$\mathbb{E}[\hat{e}(\xi)\cdot\overline{\hat{e}(\eta)}]=\mathbb{E}[m^{3}\cdot\hat{a}(\xi)\cdot\hat{b}(\xi)\cdot\overline{\hat{a}(\eta)\cdot\hat{b}(\eta)}]=m^{3}\cdot C_{a}(\delta)\cdot C_{b}(\delta).$
Define
$S(\delta)=\left|K^{c}\cap(K^{c}-\delta)\right|\leq\left|K^{c}\right|=(m-r)^{3}.$
Note that $S(\delta)$ is exactly the number of pairs $\xi,\eta\in K^{c}$ such that $\xi-\eta=\delta$. Thus we can change the summation order:
$\mathcal{E}=\sum_{x\in X_{1}\cup X_{3}}\sum_{\delta\in\mathbb{Z}_{m}^{3}}S(\delta)\cdot C_{a}(\delta)\cdot C_{b}(\delta)\cdot\omega^{x\cdot\delta}=\sum_{\delta\in\mathbb{Z}_{m}^{3}}S(\delta)\cdot C_{a}(\delta)\cdot C_{b}(\delta)\cdot\left(\sum_{x\in X_{1}\cup X_{3}}\omega^{x\cdot\delta}\right).$
Define $\mathcal{E}_{i}=\sum_{\delta,\left\lVert\delta\right\rVert_{0}=i}S(\delta)C_{a}(\delta)C_{b}(\delta)T(\delta)$, for $i\in\left\{0,\ldots,3\right\}$. Denote $T(\delta)=\sum_{x\in X_{1}\cup X_{3}}\omega^{x\cdot\delta}$. The next lemma gives us estimates on this quantity. The proof is deferred to Appendix D.
Lemma 9.8.
The following estimate holds:
$\left|T(\delta)\right|\leq\begin{cases}2m^{2}&\delta\in\left\{0,(0,*,0),(0,0,*)\right\},\\ 2m&\delta\in\left\{(*,0,0)\right\}\text{ or }\left\lVert\delta\right\rVert_{0}=2,\\ 2&\left\lVert\delta\right\rVert_{0}=3.\end{cases}$
Corollary 9.9.
It holds that $\mathcal{E}=\sum_{i=0}^{3}\mathcal{E}_{i}=O\left({(m-r)^{3}\cdot\sigma_{A}^{2}\cdot\sigma_{B}^{2}}\right)$.
Proof of 9.9.
Using the estimates we obtained in 9.7 and 9.8, we obtain the following table, which contains upper bounds on each quantity. For readability, we omit the $\frac{4\mu_{A}^{2}}{m^{3}}$ factor from the estimates (it is a lower order term that doesn’t change the result). Moreover, when writing (,0,0)(*,0,0) for example, we mean that $\delta_{1}\not=0$ while $\delta_{2}=\delta_{3}=0$.
$\delta$ Count $S(\delta)$ $\left|C_{a}(\delta)\right|$ $\left|C_{b}(\delta)\right|$ $\left|T(\delta)\right|$ $\textbf{Contribution to }\mathcal{E}$
0 11 $(m-r)^{3}$ $2\sigma_{A}^{2}/m$ $2\sigma_{B}^{2}/m$ $2m^{2}$ $(m-r)^{3}\cdot 8\sigma_{A}^{2}\sigma_{B}^{2}$
(,0,0)(*,0,0) (m1)(m-1) $(m-r)^{3}$ $2\sigma_{A}^{2}/m$ $2\sigma_{B}^{2}/m$ $2m$ $(m-r)^{3}\cdot 8\sigma_{A}^{2}\sigma_{B}^{2}\cdot\frac{m-1}{m}$
(0,,0)(0,*,0) (m1)(m-1) $(m-r)^{3}$ $2\sigma_{A}^{2}/m^{2}$ $2\sigma_{B}^{2}/m$ $2m^{2}$ $(m-r)^{3}\cdot 8\sigma_{A}^{2}\sigma_{B}^{2}\cdot\frac{m-1}{m}$
(0,0,)(0,0,*) (m1)(m-1) $(m-r)^{3}$ $2\sigma_{A}^{2}/m$ $2\sigma_{B}^{2}/m^{2}$ $2m^{2}$ $(m-r)^{3}\cdot 8\sigma_{A}^{2}\sigma_{B}^{2}\cdot\frac{m-1}{m}$
$\left\lVert\delta\right\rVert_{0}=2$ $3(m-1)^{2}$ $(m-r)^{3}$ $2\sigma_{A}^{2}/m^{2}$ $2\sigma_{B}^{2}/m^{2}$ $2m$ $\frac{(m-r)^{3}}{m}\cdot 3\cdot 8\sigma_{A}^{2}\sigma_{B}^{2}\cdot\left(\frac{m-1}{m}\right)^{2}$
$\left\lVert\delta\right\rVert_{0}=3$ $(m-1)^{3}$ $(m-r)^{3}$ $2\sigma_{A}^{2}/m^{3}$ $\sigma_{B}^{2}/m^{3} 22$ $\frac{(m-r)^{3}}{m^{3}}\cdot 8\sigma_{A}^{2}\sigma_{B}^{2}\cdot\left(\frac{m-1}{m}\right)^{3}$
Using the computations done in the table above we conclude that
$\mathcal{E}=O((m-r)^{3}\cdot\sigma_{A}^{2}\cdot\sigma_{B}^{2}),$
where the constant is hiding in the $O$-notation is around $\approx 32$ (depending how big $m$ is). ∎
We will also need the following concentration bound, which is obtained by applying sub-Gaussian concentration bounds. The proof is deferred to Appendix D.
Lemma 9.10.
In the conditions and notations of Theorem 9.6, the random variable $Y=\left\lVert a\right\rVert_{2}^{2}\cdot\left\lVert b\right\rVert_{2}^{2}$ satisfies $\mathbb{E}[Y]=\nu_{A}\nu_{B}\cdot 4(m-1)^{4}$ and $\Pr(Y<\frac{1}{2}\mathbb{E}[Y])\leq 4\exp(-\gamma(m-1)^{2})$.
We are now ready to prove the theorem:
Proof of Theorem 9.6.
First we note that $\sum_{x\in X_{1}\cup X_{3}}\left|e(x)\right|^{2}=\left\lVert I_{C}^{\top}(c^{\prime}-a*b)\right\rVert_{2}^{2}$, by definition of the STPP sets. Define $X=\sum_{x\in X_{1}\cup X_{3}}\left|e(x)\right|^{2}$ and $Y=\left\lVert a\right\rVert_{2}^{2}\cdot\left\lVert b\right\rVert_{2}^{2}$. Our goal is to bound $\mathbb{E}[X/Y]$. The first step is to reduce this to a bound on $\mathbb{E}[X]/\mathbb{E}[Y]$, which is easier handle, since $\mathbb{E}[Y]=4(m-1)^{4}\cdot\nu_{A}\cdot\nu_{B}$ and $\mathbb{E}[X]=\mathcal{E}$.
By 9.10, with probability at most $4\exp(-\gamma(m-1)^{2})$ we have $Y<\frac{1}{2}\mathbb{E}[Y]$, hence
$\mathbb{E}[X/Y]\leq\mathbb{E}\left[\frac{X}{Y}\mid Y<\frac{1}{2}\mathbb{E}[Y]\right]\cdot 4\exp(-\gamma(m-1)^{2})+\frac{2\mathbb{E}[X]}{\mathbb{E}[Y]}\cdot 1.$
Since $I_{C}^{\top}$ is a contraction, we have $X\leq\left\lVert c^{\prime}-a*b\right\rVert_{2}^{2}=\left\lVert e\right\rVert_{2}^{2}=\|{\widehat{a*b}|_{K^{c}}}\|_{2}^{2}$, where we used the linearity of the Fourier transform and Parseval. Since the projection onto $K^{c}$ is also a contraction, we have $X\leq\|\widehat{a*b}\|_{2}^{2}=\left\lVert a*b\right\rVert_{2}^{2}$ which by Cauchy-Schwarz and Young’s inequality satisfies $\left\lVert a*b\right\rVert_{2}^{2}\leq 2(m-1)^{2}\cdot\left\lVert a\right\rVert_{2}^{2}\cdot\left\lVert b\right\rVert_{2}^{2}$. Therefore $X/Y\leq 2(m-1)^{2}$ always, concluding
$\mathbb{E}[X/Y]\leq 8(m-1)^{2}\cdot\exp(-\gamma(m-1)^{2})+\frac{2\mathbb{E}[X]}{\mathbb{E}[Y]},$
so we have reduced the problem to the ratio of expectations.
By 9.9 we obtain that
$\frac{2\mathbb{E}[X]}{\mathbb{E}[Y]}=\frac{\mathcal{E}}{2(m-1)^{4}\cdot\nu_{A}\cdot\nu_{B}}=\frac{O((m-r)^{3}\cdot\sigma_{A}^{2}\cdot\sigma_{B}^{2})}{2(m-1)^{4}\cdot\nu_{A}\cdot\nu_{B}}=O\left(\frac{(m-r)^{3}\cdot\sigma_{A}^{2}\cdot\sigma_{B}^{2}}{(m-1)^{4}\cdot\nu_{A}\cdot\nu_{B}}\right).$
The hidden constant is in the vicinity of $16$ (again, depending on the size of $m$, but for $m$ which is not too small, this holds). ∎
The added exponential factor is what we pay for choosing this type of analysis, but it is negligible for large enough $m$ (assuming the distribution is “nice” enough).
9.2.3 Applications of the Theorem
An immediate corollary is the application of the theorem for the Rademacher and Gaussian distributions.
Corollary 9.11.
Given random Rademacher or Gaussian matrices $\mathbf{A}_{1},\mathbf{A}_{2},\mathbf{B}_{1},\mathbf{B}_{2}$ of size $(m-1)\times(m-1)$, and a parameter $r\leq m$, computing the set of frequencies $K$ defined in Theorem 9.6 results in matrices $\mathbf{C}_{1},\mathbf{C}_{2}$ for which
$\left\lVert\mathbf{A}_{1}\mathbf{B}_{1}-\mathbf{C}_{1}\right\rVert_{F}^{2}+\left\lVert\mathbf{A}_{2}\mathbf{B}_{2}-\mathbf{C}_{2}\right\rVert_{F}^{2}\leq O\left(\frac{(m-r)^{3}}{(m-1)^{4}}\right)\left(\left\lVert\mathbf{A}_{1}\right\rVert_{F}^{2}+\left\lVert\mathbf{A}_{2}\right\rVert_{F}^{2}\right)\left(\left\lVert\mathbf{B}_{1}\right\rVert_{F}^{2}+\left\lVert\mathbf{B}_{2}\right\rVert_{F}^{2}\right),$
in expectation.
Proof.
In both cases, we have $\mu=0$, $\nu=1$, $\sigma=1$ and $p^{2}=1$. The normalized error is thus expected to be at most
$C\cdot\left(\frac{(m-r)^{3}}{(m-1)^{4}}\right)+8\cdot(m-1)^{2}\cdot\exp\left((6-4\sqrt{2})\cdot(m-1)^{2}\right),$
where $C\approx 16$ is a global constant. ∎
We also note that $K$ has nice computation properties:
Lemma 9.12.
Let $K\subset\mathbb{Z}_{m}^{3}$ be as described in Theorem 9.6. Then:
  1. 1.
    Computing the partial Fourier transforms $\hat{a}|_{K}$ and $\hat{b}|_{K}$ can be done in $O(rm^{2})$ time.
  2. 2.
    Given a signal $f\in\mathbb{C}[\mathbb{Z}_{m}^{3}]$ such that $\hat{f}$ is supported on $K$, and given $U\subset\mathbb{Z}_{m}^{3}$, the partial inverse Fourier transform $\mathcal{F}^{-1}(\hat{f})|_{U}$ on $U$ can be computed in $O(rm^{2}\log m+\left|U\right|\cdot r)$ time.
The proof of the lemma is deferred to Appendix D. We can now deduce a general result for square matrices:
Corollary 9.13.
Given random Rademacher or Gaussian matrices $\mathbf{A},\mathbf{B}$ of size $n\times n$ where $n$ is an even number, let $m=\frac{n}{2}+1$ and let $r\leq m$ be a given parameter. Then we can compute a matrix $\mathbf{C}$ satisfying
$\left\lVert\mathbf{C}-\mathbf{A}\mathbf{B}\right\rVert_{F}^{2}\leq O\left(\frac{\left(\frac{n}{2}+1-r\right)^{3}}{\left(\frac{n}{2}\right)^{4}}\right)\cdot\left\lVert\mathbf{A}\right\rVert_{F}^{2}\cdot\left\lVert\mathbf{B}\right\rVert_{F}^{2}=O\left(\frac{(n+2-2r)^{3}}{n^{4}}\right)\cdot\left\lVert\mathbf{A}\right\rVert_{F}^{2}\cdot\left\lVert\mathbf{B}\right\rVert_{F}^{2},$
in expectation, using the Fourier truncation approach. The matrix $\mathbf{C}$ can be computed in time $O(rn^{2})$.
Proof.
Combine 9.11 with D.1 to obtain the error bound. For the runtime, this is exactly 9.12. ∎
10 Empiric Trials
We will present results for two empiric trials, that compare the performance of different algorithms for different values of $r\leq n$, where $n$ is the size of the randomly drawn matrices. The algorithms are:
  1. 1.
    JL Sketch: Based on [Woo14] and 7.1, given $\mathbf{A},\mathbf{B}$ we create a matrix $S\in\mathbb{R}^{n\times r}$ by randomly drawing elements out of $\mathcal{N}(0,\frac{1}{\sqrt{r}})$. We then return $(\mathbf{A}S)(S^{\top}\mathbf{B})$.
  2. 2.
    PolyForm: Based on Theorem 8.7, we embed $\mathbf{A},\mathbf{B}$ into $\mathbb{Z}_{rn^{2}}$ using approximate TPP sets based on arithmetic progressions. We then compute the convolution and return a matrix based on a subset of the coordinates of the convolution.
  3. 3.
    TPP Fourier Truncation: Based on the Fourier domain truncation approach, we embed $\mathbf{A},\mathbf{B}$ into $\mathbb{Z}_{n^{3}}$ using the Vanilla TPP construction, and then compute $\hat{c}|_{K}$, where $K$ is the set of frequencies
    $K=\left\{\xi<\frac{rn^{2}}{2}\right\}\cup\left\{\xi\equiv 0\pmod{n^{2}}\right\}\cup\left\{n^{3}-\frac{rn^{2}}{2}\leq\xi\right\}.$
    Note that $\left|K\right|\leq rn^{2}+n$. While we do not know of a way to compute this partial transform in time $\tilde{O}(rn^{2})$ (in fact, the inverse transform given $\hat{c}|_{K}$ can be computed in $\tilde{O}(rn^{2})$ using an algorithm proposed by [DHR97]), we did identify this set to be the heavy set, due to the structure of the underlying support sets (a similar argument to the one carried out in Theorem 9.6).
  4. 4.
    STPP Fourier Truncation: Based on Theorem 9.6, we split the matrices $\mathbf{A},\mathbf{B}$ into four blocks, and invoke the Fourier truncation algorithms $4$ times, each time to compute two block products, with $m=\frac{n}{2}+1$. We truncate the Fourier domain to the slabs proposed in the theorem, but instead of the slabs having width $r$ (which would result in retaining $\leq 3rm^{2}\approx rn^{2}$ frequencies).
  5. 5.
    $r$-$\mathsf{SVD}$: taking the top $r$ singular values to obtain the best rank $r$-approximation.
We run the trial with $10$ different scales of $r$, taking linear jumps between $1$ and $n$ (or $m$ in the STPP case). For each value of $r$ we run $10$ different trials and present the results as Violin Plots, which aim to show how the different trials have been distributed. The comparison can be understood as – how well does each algorithm perform given the budget to compute $\approx rn^{2}$ multiplications (in the (Normal Form)., as depicted in the formulation of 6.1). The red line in the graphs is the reference SOTA ratio $1/r$.
Some of the plots include a matching plot for the observed ranks of the output matrices. Computing the nuclear norm (i.e., sum of squared singular values) of the output matrices instead of the rank reveals that the full rank observed in our methods is not degenerate (i.e., caused by very small non-zero singular values). In fact it strengthens our empirical observation that our methods produce (effective) high rank outputs.
Figure 3: Comparison of item 1, item 2 and item 3 on random $\left\{\pm 1\right\}$ matrices of size $200\times 200$.
Figure 4: Comparison of item 3 and item 4 on random $\left\{0,1\right\}$ matrices of size $200\times 200$.
Figure 5: Comparison of item 4 and item 5 on random $\left\{\pm 1\right\}$ matrices of size $400\times 400$.
It is notable that the error on $\left\{\pm 1\right\}$ matrices are $\approx$ twice the error on $\left\{0,1\right\}$, and that the errors are highly concentrated (the violins are “squashed”), which fits our results.
11 Conclusions & Open Questions
The paper presents very simple and practical algorithms for approximate MM that potentially have better qualities compared with standard sketching techniques. It provides a theoretical foundation and tools to analyze the theoretical guarantees of the algorithms. A few open questions that we would like to note:
  • Is there a simpler, more elegant proof for Theorem 1.4? Can this result can be generalized for larger classes of distributions over matrices? In particular, the empiric evidence suggest that the same approach in the TPP case yields good results too.
  • What is the relationship between input and output rank of the matrices in the algorithm in section 8? We see empiric evidence (see Figure 3) that the algorithm produces high rank outputs.
  • Studying the (Approximate Triple Product Quantity). in the case of non-abelian groups. Can we prove general lower bounds?
  • Finding approximate STPP constructions (in the sense of (Approximate Triple Product Quantity).). We believe such constructions will lead to new state-of-the-art trade-off in the worst case, using similar ideas to section 8.
  • Are there other low-degree polynomial representations that might be useful? For example, using the theory of Chebyshev approximation, which deals with uniform approximation of polynomials over compact domains, sounds useful. In our attempt, it did not yield good results.
  • On the practical side, it is interesting to test this framework on real life matrices, specifically in the context of DL.
Acknowledgments
We thank Zachary Chase for insightful discussions on the PolyForm scheme and for originating the proof of the lower bound (8.5).
Appendix A Fast Combinatorial Algorithm
For convenience we introduce the following notation - for $i\in\left\{1,2,3\right\}$ and $\boldsymbol{\alpha}\in\left\{0,1\right\}^{N}$ we let $S_{i}^{\boldsymbol{\alpha}}$ denote the (cartesian) product $S_{i}^{\boldsymbol{\alpha}}:=\prod_{j=1}^{N}S_{i}^{\boldsymbol{\alpha}_{j}}$. We also let $\mathsf{bin}\left(x\right)\in\left\{0,1\right\}^{N}$ denote the binary representation of a number $x\in\left[2^{N}\right]$ (for convenience we identify 0 with $2^{N}$).
Algorithm 1 $\mathbb{Z}_{m}^{3N}$ STPP algorithm ($m,N$)
1:Matrices $\mathbf{A}_{1},\ldots,\mathbf{A}_{2^{N}},\mathbf{B}_{1},\ldots,\mathbf{B}_{2^{N}}$ of size $\left(m-1\right)^{N}\times\left(m-1\right)^{N}$.
2:Product matrices $\mathbf{A}_{1}\mathbf{B}_{1},\ldots,\mathbf{A}_{2^{N}}\mathbf{B}_{2^{N}}$.
3:for $x=1,\ldots,2^{N}$ do
4:  Index $\mathbf{A}_{i}$ by $S_{2}^{\mathsf{bin}(x)}-S_{1}^{\mathsf{bin}(x)}\triangleright$ Difference Set
5:end for
6:Compute the FFT of the resulting signal $a$ and denote the result by $\hat{a}$
7:for $x=1,\ldots,2^{N}$ do
8:  Index $\mathbf{B}_{i}$ by $S_{3}^{\mathsf{bin}(x)}-S_{2}^{\mathsf{bin}(x)}$
9:end for
10:Compute the FFT of the resulting signal $b$ and denote the result by $\hat{b}$
11:Compute the element-wise product $\hat{c}=\hat{a}\odot\hat{b}$
12:Compute the inverse FFT of $\hat{c}$, denoted by $c$
13:for $x=1,\ldots,2^{N}$ do
14:  Read $\mathbf{A}_{x}\mathbf{B}_{x}$ from the coefficients of $S_{3}^{\mathsf{bin}(x)}-S_{1}^{\mathsf{bin}(x)}$
15:end for
16:return $\mathbf{A}_{1}\mathbf{B}_{1},\ldots,\mathbf{A}_{2^{n}}\mathbf{B}_{2^{N}}$
Appendix B More Definitions
For completeness we present here a few definitions used in the paper.
Definition B.1 (Normal Form).
Let $\mathbb{F}$ denote a field and let $f:\mathbb{F}^{n}\times\mathbb{F}^{m}\to\mathbb{F}^{p}$ denote a bilinear map. Then $f$ can be represented in Normal Form, via three matrices $\mathbf{U}\in\mathbb{F}^{n\times r},\mathbf{V}\in\mathbb{F}^{m\times r}$ and $\mathbf{W}\in\mathbb{F}^{r\times p}$, such that for every $\mathbf{a}\in\mathbb{F}^{n},\mathbf{b}\in\mathbb{F}^{m}$ it holds
$f(\mathbf{a},\mathbf{b})=\mathbf{W}^{\top}(\mathbf{U}\mathbf{a}\odot\mathbf{V}\mathbf{b}),$
where $\odot$ denotes the element-wise multiplication of vectors. The quantity $r$ is called the rank of this representation.
Furthermore, a trilinear tensor $\mathcal{T}$ is a formal sum over three sets of variables $\left\{X_{i}\right\}_{i\in[n]}$, $\left\{Y_{i}\right\}_{i\in[m]}$, $\left\{Z_{i}\right\}_{i\in[p]}$, given as $\sum_{i,j,k}T(i,j,k)\cdot X_{i}Y_{j}Z_{k}$, where $T:[n]\times[m]\times[p]\to\mathbb{F}$ is a function. Then $\mathcal{T}$ represents a bilinear map $f_{\mathcal{T}}:\mathbb{F}^{n}\times\mathbb{F}^{m}\to\mathbb{F}^{p}$ defined by
$f_{\mathcal{T}}(\mathbf{a},\mathbf{b})]_{k}:=\sum_{i,j}T(i,j,k)\cdot\mathbf{a}_{i}\cdot\mathbf{b}_{j}.$
The normal form shows that the converse is also true, i.e., given a bilinear form and a representation $\mathbf{U},\mathbf{V},\mathbf{W}$ as above, we set
$\mathcal{T}=\sum_{t=1}^{r}\left(\sum_{i}\mathbf{U}_{t,i}X_{i}\right)\left(\sum_{j}\mathbf{V}_{t,j}Y_{j}\right)\left(\sum_{k}\mathbf{W}_{k,t}Z_{k}\right),$
and obtain $f_{\mathcal{T}}=f$. The tensor rank of a tensor is the minimal number $r$ for which $\mathcal{T}$ can be represented as the sum of simple tensors, which are products as in the previous expression. Thus, the rank of the representation is an upper bound on the tensor rank. We denote the tensor rank by $\mathsf{Trk}(\mathcal{T})$.
For completeness, we present the definition of the simultaneous triple product property (STPP) capturing when the group structure tensor $\mathcal{T}_{G}$ (defined via $T_{G}(g,h,f)=1$ iff $gh=f$, where $g,h,f\in G$ are our variables) can be restricted to a direct sum of matrix multiplication tensors. This definition was given in [Coh+05].
Definition B.2 (STPP).
We say that a finite group $G$ satisfies the STPP for $\left\langle{n_{1},m_{1},p_{1}}\right\rangle,\ldots\left\langle{n_{\ell},m_{\ell},p_{\ell}}\right\rangle$, if there exist triplets of subsets in $G$, $\left\{(S_{i},T_{i},U_{i})\right\}_{i=1,\ldots,\ell}$, of sizes $\left|S_{i}\right|=n_{i},\left|T_{i}\right|=m_{i},\left|U_{i}\right|=p_{i}$, such that $(S_{i},T_{i},U_{i})$ satisfy the TPP, and they don’t mix:
$\forall i,j,k\in\left\{1,\ldots,\ell\right\}:\quad s_{i}\cdot(s_{j}^{\prime})^{-1}\cdot t_{j}\cdot(t_{k}^{\prime})^{-1}\cdot u_{k}\cdot(u_{i}^{\prime})^{-1}=1_{G}\implies i=j=k,$
assuming $x_{a},x_{a}^{\prime}\in X_{a}$ (with $X\in\left\{S,T,U\right\},a\in\left\{i,j,k\right\}$).
An important and useful property is:
Lemma B.3.
If $\left\langle{S,T,U}\right\rangle$ satisfy the TPP in $G$ and $\left\langle{S^{\prime},T^{\prime},U^{\prime}}\right\rangle$ in $G^{\prime}$, then $\left\langle{S\times S^{\prime},T\times T^{\prime},U\times U^{\prime}}\right\rangle$ satisfies the TPP in the product group $G\times G^{\prime}$. Similarly, if $\left\{\left\langle{S_{i},T_{i},U_{i}}\right\rangle\right\}_{i=1}^{k}$ satisfy the STPP in $G$ and $\left\{\left\langle{S_{i}^{\prime},T_{i}^{\prime},U_{i}^{\prime}}\right\rangle\right\}_{i=1}^{k^{\prime}}$ in $G^{\prime}$, then $\left\{\left\langle{S_{i}\times S_{j}^{\prime},T_{i}\times T_{j}^{\prime},U_{i}\times U_{j}^{\prime}}\right\rangle\right\}_{i\in[k],j\in[k^{\prime}]}$ satisfy STPP in $G\times G^{\prime}$.
Appendix C Low Degree Embedding Proofs
In this section we give the proof of Theorem 8.6 and Theorem 8.7. We first introduce some notation.
Definition C.1 (Collisions).
Let $G$ denote a finite group, $\left|G\right|\geq n^{2}$ and $(S,T,U)\in\mathcal{I}(G,n)$ and indexing triplet. We define the set of Collisions with respect to $G,n$ and (Indexing Triplet). $(S,T,U)$ by:
$\mathrm{Collisions}_{G}\left(S,T,U\right):=\left\{(s,t,t^{\prime},u)\mid t\not=t^{\prime},\ \exists x\in S,y\in U:\ s^{-1}t{t^{\prime}}^{-1}u=x^{-1}y\right\}.$
In words, elements of $\mathrm{Collisions}_{G}\left(S,T,U\right)$ correspond with coefficients that would be wrongfully mapped into an index of interest (i.e. used to index an element of $\mathbf{A}\mathbf{B}$).
Definition C.2 (Collisions & Buckets).
In the same conditions as above, we define the set of Collisions for an index $x,y$, where $x\in S,y\in U$ by:
$\mathrm{Collisions}_{G}^{x,y}\left(S,T,U\right)=\left\{(s,t,t^{\prime},u)\mid t\not=t^{\prime},s^{-1}t{t^{\prime}}^{-1}u=x^{-1}y\right\}.$
We also define the Bucket of $x,y$ by:
$\mathrm{Bucket}_{G}^{x,y}\left(S,T,U\right)=\left\{(s,t,t^{\prime},u)\mid s^{-1}t{t^{\prime}}^{-1}u=x^{-1}y\right\}.$
We note the following useful characterization of the error.
Lemma C.3.
Let $G$ denote a finite group, $\left|G\right|\geq n^{2}$, $\mathbf{A},\mathbf{B}\in\mathbb{R}^{n}$ be two given matrices and $(S,T,U)\in\mathcal{I}(G,n)$ some (Indexing Triplet).. Let $\mathbf{C}$ denote the output of the deterministic PolyForm algorithm. Then for every $x\in S,y\in U$ we have
$\mathbf{C}_{x,y}=\sum_{(s,t,t^{\prime},u)\in\mathrm{Bucket}_{G}^{x,y}\left(S,T,U\right)}\mathbf{A}_{s,t}\mathbf{B}_{t^{\prime}u},$
and the error is
$\left\lVert\mathbf{C}-\mathbf{A}\mathbf{B}\right\rVert_{F}^{2}=\sum_{x\in S,y\in U}\left(\sum_{(s,t,t^{\prime},u)\in\mathrm{Collisions}_{G}^{x,y}\left(S,T,U\right)}\mathbf{A}_{s,t}\mathbf{B}_{t^{\prime},u}\right)^{2}.$
Proof.
Follows from opening up the definition of a group algebra product, noting that
$(\mathbf{A}\mathbf{B})_{x,y}=\sum_{\begin{subarray}{c}s\in S,t\in T,u\in U\\ s^{-1}(tt^{-1})u=x^{-1}y\end{subarray}}\mathbf{A}_{s,t}\mathbf{B}_{t,u},$
leaving us with a sum over $(s,t,t^{\prime},u)\in\mathrm{Collisions}_{G}^{x,y}\left(S,T,U\right)$ in the coordinate $x,y$ of $\mathbf{C}-\mathbf{A}\mathbf{B}$. ∎
Proof of Theorem 8.6.
For brevity, we let $\mathcal{R}$ denote a randomized PolyForm setup, composed of an indexing triplet $(S,T,U)\in\mathcal{I}(G,n)$, permutations of the indexing sets and sign vectors $\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{\gamma}$. We denote the deterministic error from this setup by $\mathcal{E}_{\mathcal{R}}(\mathbf{A},\mathbf{B})=\frac{\left\lVert\mathbf{C}-\mathbf{A}\mathbf{B}\right\rVert_{F}^{2}}{\left\lVert\mathbf{A}\right\rVert_{F}^{2}\cdot\left\lVert\mathbf{B}\right\rVert_{F}^{2}}$, where $\mathbf{C}$ is the output matrix of the algorithm with setup $\mathcal{R}$. By Yao’s principle, for any distribution $\mathcal{D}$ over inputs of the algorithm (i.e. couples of matrices), it holds
$\mathcal{E}_{G}\geq\min_{\mathcal{R}}\underset{(\mathbf{A},\mathbf{B})\sim\mathcal{D}}{\mathbb{E}}\left[\mathcal{E}_{\mathcal{R}}(\mathbf{A},\mathbf{B})\right].$
Set $\mathcal{D}$ to be the uniform distribution on Gaussian matrices (all entries are independently drawn uniformly from $\{\pm 1\}$). We can assume without loss of generality that the permutations are all the identity permutations, because $\mathcal{D}$ is invariant under permutation. Moreover, $\left\lVert\mathbf{A}\right\rVert_{F}^{2}=\left\lVert\mathbf{B}\right\rVert_{F}^{2}=n^{2}$ when $(A,B)\sim\mathcal{D}$. Thus, by C.3 and the fact that $(\mathbf{A}\odot\boldsymbol{\alpha}\boldsymbol{\gamma}^{\top})_{s,t}=\boldsymbol{\alpha}_{s}\boldsymbol{\gamma}_{t}\mathbf{A}_{s,t}$ (and similarly for $\mathbf{B}\odot\boldsymbol{\gamma}\boldsymbol{\beta}^{\top}$), we have
$n^{4}\cdot\underset{(\mathbf{A},\mathbf{B})\sim\mathcal{D}}{\mathbb{E}}\left[\mathcal{E}_{\mathcal{R}}(\mathbf{A},\mathbf{B})\right]$ $=\underset{}{\mathbb{E}}\left[\sum_{x\in S,y\in U}\left(\sum_{(s,t,t^{\prime},u)\in\mathrm{Collisions}_{G}^{x,y}\left(\mathcal{R}\right)}\boldsymbol{\alpha}_{s}\boldsymbol{\gamma}_{t}\boldsymbol{\gamma}_{t^{\prime}}\boldsymbol{\beta}_{u}\cdot\mathbf{A}_{s,t}\mathbf{B}_{t^{\prime},u}\right)^{2}\right]$
$=\sum_{x\in S,y\in U}\underset{}{\mathbb{E}}\left[\left(\sum_{(s,t,t^{\prime},u)\in\mathrm{Collisions}_{G}^{x,y}\left(\mathcal{R}\right)}\boldsymbol{\alpha}_{s}\boldsymbol{\gamma}_{t}\boldsymbol{\gamma}_{t^{\prime}}\boldsymbol{\beta}_{u}\cdot\mathbf{A}_{s,t}\mathbf{B}_{t^{\prime},u}\right)^{2}\right].$ (1)
Fixing $x\in S,y\in U$, we note that
$\underset{}{\mathbb{E}}\left[\left(\sum_{(s,t,t^{\prime},u)\in\mathrm{Collisions}_{G}^{x,y}\left(\mathcal{R}\right)}\boldsymbol{\alpha}_{s}\boldsymbol{\gamma}_{t}\boldsymbol{\gamma}_{t^{\prime}}\boldsymbol{\beta}_{u}\cdot\mathbf{A}_{s,t}\mathbf{B}_{t^{\prime},u}\right)^{2}\right]$
$=\sum_{(s,t,t^{\prime},u)\in\mathrm{Collisions}_{G}^{x,y}\left(\mathcal{R}\right)}\mathbb{E}\left[\left(\boldsymbol{\alpha}_{s}\boldsymbol{\gamma}_{t}\boldsymbol{\gamma}_{t^{\prime}}\boldsymbol{\beta}_{u}\cdot\mathbf{A}_{s,t}\mathbf{B}_{t^{\prime},u}\right)^{2}\right]$ (2)
$+\sum_{\begin{subarray}{c}(s,t,t^{\prime},u)\\ \not=\\ (p,r,r^{\prime},q)\end{subarray}\in\mathrm{Collisions}_{G}^{x,y}\left(\mathcal{R}\right)}\mathbb{E}\left[\left(\boldsymbol{\alpha}_{s}\boldsymbol{\gamma}_{t}\boldsymbol{\gamma}_{t^{\prime}}\boldsymbol{\beta}_{u}\cdot\mathbf{A}_{s,t}\mathbf{B}_{t^{\prime},u}\right)\cdot\left(\boldsymbol{\alpha}_{p}\boldsymbol{\gamma}_{r}\boldsymbol{\gamma}_{r^{\prime}}\boldsymbol{\beta}_{q}\cdot\mathbf{A}_{p,r}\mathbf{B}_{r^{\prime},q}\right)\right].$ (3)
Since $\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{\gamma}$ and $\mathbf{A},\mathbf{B}$ all take values in $\pm 1$, we obtain
$(2)=\sum_{(s,t,t^{\prime},u)\in\mathrm{Collisions}_{G}^{x,y}\left(\mathcal{R}\right)}\mathbb{E}\left[1\right]=\left|\mathrm{Collisions}_{G}^{x,y}\left(\mathcal{R}\right)\right|.$
Regarding $(3)$, the assumption $(s,t,t^{\prime},u)\not=(p,r,r^{\prime},q)$ implies that $(s,t)\not=(p,r)$ or $(t^{\prime},u)\not=(r^{\prime},q)$, hence at most one couple of random variables in the set $\{\mathbf{A}_{s,t},\mathbf{A}_{p,r},\mathbf{B}_{t^{\prime},u},\mathbf{B}_{r^{\prime},q}\}$ are not independent. Using the multiplicity of the expectation, we obtain
$(3)=\sum_{\begin{subarray}{c}(s,t,t^{\prime},u)\\ \not=\\ (p,r,r^{\prime},q)\end{subarray}\in\mathrm{Collisions}_{G}^{x,y}\left(\mathcal{R}\right)}\mathbb{E}\left[\boldsymbol{\alpha}_{s}\boldsymbol{\gamma}_{t}\boldsymbol{\gamma}_{t^{\prime}}\boldsymbol{\beta}_{u}\cdot\boldsymbol{\alpha}_{p}\boldsymbol{\gamma}_{r}\boldsymbol{\gamma}_{r^{\prime}}\boldsymbol{\beta}_{q}\right]\cdot\mathbb{E}\left[\mathbf{A}_{s,t}\mathbf{B}_{t^{\prime},u}\cdot\mathbf{A}_{p,r}\mathbf{B}_{r^{\prime},q}\right]=0,$
where we used the fact that a multiple of independent Rademacher RVs ($\pm 1$ with equal probability) is 0 in expectation.
To sum up, we have shown
$(1)=\sum_{x\in S,y\in U}\left|\mathrm{Collisions}_{G}^{x,y}\left(\mathcal{R}\right)\right|=\left|\mathrm{Collisions}_{G}\left(\mathcal{R}\right)\right|\geq\rho_{n}(G)-n^{3}.$
Since $\mathcal{R}$ was arbitrary, we have thus shown that
$\mathcal{E}_{G}\geq\min_{\mathcal{R}}\underset{(\mathbf{A},\mathbf{B})\sim\mathcal{D}}{\mathbb{E}}\left[\mathcal{E}_{\mathcal{R}(\mathbf{A},\mathbf{B})}\right]\geq\frac{\rho_{n}(G)-n^{3}}{n^{4}}.$
Remark C.4.
The random signs can be drawn from a non-uniform $\left\{\pm 1\right\}$ distribution, and the proof would still work. So we can allow a slightly wider framework.
Proof of Theorem 8.7.
For the sake of clarity, we give a complete proof and explicit construction of the output matrix $\mathbf{C}$. Let $\mathbf{A},\mathbf{B}$ denote two given square matrices, denote $D=rn^{2}$. The proof has two parts - construction of $\mathbf{C}$ and analysis of the error.
Construction. By definition, there exists $S,T,U\subset\mathbb{Z}_{D}$, an indexing triplet, for which $\rho_{n,G}(S,T,U)=\rho_{n}(G)$. Define some random order over the elements of $S,T,U$, denoted as $\left(X\left(i\right)\right)_{i=0}^{n-1}$ (for $X=S,T,U$). Denote $\mathbf{A}=\left(\mathbf{A}_{i,k}\right)_{i,k\in\left[n\right]}$ and $\mathbf{B}=\left(\mathbf{B}_{l,j}\right)_{l,j\in\left[n\right]}$. At last, draw three independent sign vectors $\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{\gamma}:\left[n\right]\to\left\{\pm 1\right\}$ with uniform distribution.
Define the following polynomials -
$P_{A}\left(x\right)$ $=\sum_{i,k\in\left[n\right]}\boldsymbol{\alpha}_{i}\cdot\boldsymbol{\gamma}_{k}\cdot\mathbf{A}_{i,k}\cdot x^{S\left(i\right)+T\left(k\right)}$
$P_{B}\left(x\right)$ $=\sum_{l,j\in\left[n\right]}\boldsymbol{\beta}_{j}\cdot\boldsymbol{\gamma}_{k}\cdot\mathbf{B}_{l,j}\cdot x^{-T\left(l\right)+U\left(j\right)}$
(the addition and subtraction are group operations, i.e. modulo $D$). Compute their convolution (using the cyclic group FFT algorithm), and denote the result by $\sum_{r\in[D]}\mathbf{c}_{r}x^{r}$. Now, define $\mathbf{C}\in\mathbb{R}^{n\times n}$ by -
$\left(\mathbf{C}\right)_{i,j}:=\left(\boldsymbol{\alpha}_{i}\cdot\boldsymbol{\beta}_{j}\right)\cdot\mathbf{c}_{i\cdot S\left(i\right)+U\left(j\right)}\qquad\left(\forall i,j\in\left[n\right]\right).$
Constructing the polynomials is done in $O\left(D\right)$ time, computing the convolution in $O\left(D\log\left(D\right)\right)$ time and constructing $\mathbf{C}$ in $O\left(n^{2}\right)$ time. Overall we reach $O\left(rn^{2}\log\left(n\right)\right)$ running time.
Analysis. Fix $i,j\in\left[n\right]$. We have by definition of $\mathbf{C}$:
$\left(\boldsymbol{\alpha}_{i}\cdot\boldsymbol{\beta}_{j}\right)\cdot\left(\mathbf{C}\right)_{i,j}$ $=\mathbf{c}_{S\left(i\right)+U\left(j\right)}$
$=\sum_{\begin{subarray}{c}i^{\prime},j^{\prime},k,k^{\prime}\in\left[n\right]\\ \left[S\left(i^{\prime}\right)+T\left(k\right)\right]+\left[U\left(j^{\prime}\right)-T\left(k^{\prime}\right)\right]\\ =S\left(i\right)+U\left(j\right)\end{subarray}}\left(\boldsymbol{\alpha}_{i^{\prime}}\cdot\boldsymbol{\gamma}_{k}\cdot\mathbf{A}_{i^{\prime},k}\right)\cdot\left(\boldsymbol{\beta}_{j^{\prime}}\cdot\boldsymbol{\gamma}_{k^{\prime}}\cdot\mathbf{B}_{k^{\prime},j^{\prime}}\right)$
We can expand and simplify this expression to obtain
$=\underbrace{\sum_{\begin{subarray}{c}i^{\prime},j^{\prime},k=k^{\prime}\in\left[n\right]\\ S\left(i^{\prime}\right)+U\left(j^{\prime}\right)=S\left(i\right)+U\left(i\right)\end{subarray}}\boldsymbol{\alpha}_{i^{\prime}}\cdot\boldsymbol{\beta}_{j^{\prime}}\cdot\boldsymbol{\gamma}_{k}^{2}\cdot\mathbf{A}_{i^{\prime},k}\mathbf{B}_{k,j^{\prime}}}_{*}+\underbrace{\sum_{\begin{subarray}{c}i^{\prime},j^{\prime},k\not=k^{\prime}\in\left[n\right]\\ S\left(i^{\prime}\right)+U\left(j^{\prime}\right)+T\left(k\right)-T\left(k^{\prime}\right)\\ =S\left(i\right)+U\left(i\right)\end{subarray}}\boldsymbol{\alpha}_{i^{\prime}}\cdot\boldsymbol{\beta}_{j^{\prime}}\cdot\boldsymbol{\gamma}_{k}\cdot\boldsymbol{\gamma}_{k^{\prime}}\cdot\mathbf{A}_{i^{\prime},k}\mathbf{B}_{k^{\prime},j^{\prime}}}_{**}.$
By definition of indexing triplets, we know that $\left|S+U\right|=n^{2}$. Thus, for the first sum *, we have that $S\left(i^{\prime}\right)+U\left(j^{\prime}\right)=S\left(i\right)+U\left(j\right)\iff\left(i^{\prime}=i\land j^{\prime}=j\right)$. This implies that we can rewrite the first sum as
$\sum_{k\in\left[n\right]}\boldsymbol{\alpha}_{i}\cdot\boldsymbol{\beta}_{j}\cdot\boldsymbol{\gamma}_{k}^{2}\cdot\mathbf{A}_{i,k}\cdot\mathbf{B}_{k,j}=(\boldsymbol{\alpha}_{i}\cdot\boldsymbol{\beta}_{j})\cdot\sum_{k\in[n]}\mathbf{A}_{i,k}\cdot\mathbf{B}_{k,j}=(\boldsymbol{\alpha}_{i}\boldsymbol{\beta}_{j})\cdot(\mathbf{A}\mathbf{B})_{i,j},,$
using the definition of matrix multiplication. This sign can be canceled by multiplying it by itself as in the definition of $\mathbf{C}$.
The second sum **, is thus the sum of "noise" terms. Formally - $\left|\mathbf{C}_{i,j}-\left(\mathbf{AB}\right)_{i,j}\right|=\left|**\right|$. As we wish to minimize the squared Frobenius norm $\left\lVert\mathbf{C}-\mathbf{A}\mathbf{B}\right\rVert_{F}^{2}$, we end up minimizing the squares of these noise terms. We further observe that -
$\forall i^{\prime},j^{\prime},k,k^{\prime}$ $\in\left[n\right],k\not=k^{\prime}:\quad S\left(i^{\prime}\right)+U\left(j^{\prime}\right)+T\left(k\right)-T\left(k^{\prime}\right)=S\left(i\right)+U\left(j\right)$
$\iff$ $T(k)-T(k^{\prime})+S(i^{\prime})-S(i)+U(j^{\prime})-U(j)=0.$
Thus when summing over all $i,j\in[n]$ we obtain that the number of items in the sum is exactly the number of non-trivial solutions to the equation defining the (Approximate Triple Product Quantity). - which is $\rho_{n}(G)-n^{3}$.
Bound. We continue to bound the error. By linearity of expectation,
$\mathbb{E}\left[\left\|\mathbf{C}-\mathbf{AB}\right\|_{F}^{2}\right]$ $=\mathbb{E}\left[\sum_{i,j}\left(\mathbf{C}_{i,j}-\left(\mathbf{AB}\right)_{i,j}\right)^{2}\right]=\sum_{i,j}\mathbb{E}\left[\left(\mathbf{C}_{i,j}-\left(\mathbf{AB}\right)_{i,j}\right)^{2}\right].$
Let $\mathbf{1}_{i,j}$ denote an indicator over quad-tuples $(i^{\prime},j^{\prime},k,k^{\prime}),k\not=k^{\prime}$ that obtains $1$ if $T(k)-T(k^{\prime})+S(i^{\prime})+U(j^{\prime})=S(i)+U(j)$. For brevity we denote such quad-tuples by $t=(i^{\prime},j^{\prime},k,k^{\prime})$ and $t^{\prime}=(p,q,l,l^{\prime})$. We denote $f_{t}=:\mathbf{1}_{i,j}(t)$ to be a random variable which relies on the ordering permutation $\pi$ applied independently on $S,T,U$. We denote $g_{t}:=\boldsymbol{\alpha}_{i^{\prime}}\boldsymbol{\beta}_{j^{\prime}}\boldsymbol{\gamma}_{k}\boldsymbol{\gamma}_{k^{\prime}}\mathbf{A}_{i^{\prime},k}\mathbf{B}_{k^{\prime},j^{\prime}}$ to be a random variable which relies on the sign vectors $\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{\gamma}$. We note that $f_{t},g_{t}$ are independent. Thus, opening up the error terms while fixing $i,j\in\left[n\right]$, we have
$\mathbb{E}\left[\left(\mathbf{C}_{i,j}-\left(\mathbf{AB}\right)_{i,j}\right)^{2}\right]$ $=\mathbb{E}\left[\left(\sum_{t}\mathbf{1}_{i,j}(t)\cdot\boldsymbol{\alpha}_{i^{\prime}}\boldsymbol{\beta}_{j^{\prime}}\boldsymbol{\gamma}_{k}\boldsymbol{\gamma}_{k^{\prime}}\cdot\mathbf{A}_{i^{\prime},k}\cdot\mathbf{B}_{k^{\prime},j^{\prime}}\right)^{2}\right]$
$=\mathbb{E}\left[\left(\sum_{t}f_{t}\cdot g_{t}\right)^{2}\right]=\mathbb{E}\left[\sum_{t,t^{\prime}}f_{t}\cdot f_{t^{\prime}}\cdot g_{t}\cdot g_{t^{\prime}}\right]=\sum_{t,t^{\prime}}\mathbb{E}\left[f_{t}\cdot f_{t^{\prime}}\right]\cdot\mathbb{E}\left[g_{t}\cdot g_{t^{\prime}}\right].$
Note that for the random signs in $g_{t}\cdot g_{t^{\prime}}$ not to cancel out in expectation, we must have that $i^{\prime}=p,j^{\prime}=q$ and $\left(k=l,k^{\prime}=l^{\prime}\right)\lor\left(k=l^{\prime},k^{\prime}=l\right)$. In all other cases we have two independent signs that cancel out in expectation. For the first case, where $t=t^{\prime}$, we note that
$\mathbb{E}\left[g_{t}\cdot g_{t^{\prime}}\right]=\mathbf{A}_{i^{\prime},k}\cdot\mathbf{B}_{k^{\prime},j^{\prime}}\cdot\mathbf{A}_{i^{\prime},k}\cdot\mathbf{B}_{k^{\prime},j^{\prime}}=\mathbf{A}_{i^{\prime},k}^{2}\cdot\mathbf{B}_{k^{\prime},j^{\prime}}^{2}\geq 0.$
For the second case, where $t^{\prime}=\left(i^{\prime},j^{\prime},k^{\prime},k\right)$, we note that $f_{t}$ and $f_{t^{\prime}}$ are fully-correlated. Assuming $f_{t}=1$ (i.e. $t$ is an error term in $\mathbf{C}_{i,j}$) then $f_{t^{\prime}}=1$. Thus
$S\left(i^{\prime}\right)+T\left(k\right)+U\left(j^{\prime}\right)-T\left(k^{\prime}\right)$ $=S\left(i^{\prime}\right)+T\left(k^{\prime}\right)+U\left(j^{\prime}\right)-T\left(k\right)$
$\iff$ $T\left(k\right)=T\left(k^{\prime}\right)$
$\iff$ $k=k^{\prime}$
which contradicts the assumption that $t\not=t^{\prime}$. In other words, in the second case, we must have $f_{t}=0$, which implies
$\mathbb{E}\left[f_{t}\cdot f_{t^{\prime}}\right]\cdot\mathbb{E}\left[g_{t}\cdot g_{t^{\prime}}\right]=0\cdot\mathbb{E}\left[g_{t}\cdot g_{t^{\prime}}\right]=0.$
To conclude the discussion, we note that each $t\in I$ can only fall into a single “bucket” (be an error term of a single index in $\mathbf{C}$). Thus the sum is disjoint for each $i,j$. Per the analysis above, our sum simplifies to:
$\sum_{i,j}\mathbb{E}\left[\left(\mathbf{C}_{i,j}-\left(\mathbf{AB}\right)_{i,j}\right)^{2}\right]$ $=\sum_{i,j}\left(\sum_{t=t^{\prime}\in I}\mathbb{E}\left[f_{t}\cdot f_{t^{\prime}}\cdot g_{t}\cdot g_{t^{\prime}}\right]\right)$
$\left(\text{disjoint sum}\right)$ $=\sum_{i,j}\sum_{k\not=k^{\prime}}\mathbb{E}\left[f_{t}^{2}\cdot g_{t}^{2}\right]$
$\left(\text{independence}\right)$ $\leq\sum_{i,j,k,k^{\prime}}\mathbb{E}\left[f_{t}^{2}\right]\cdot\mathbb{E}\left[g_{t}^{2}\right]$
$\left(\mathbb{E}\left[f_{t}^{2}\right]=\mathbb{E}\left[f_{t}\right]\text{ since }f_{t}\sim\text{Ber}\right)$ $=\frac{\rho_{n}(G)-n^{3}}{n^{4}}\cdot\sum_{i,j,k,k^{\prime}}a_{i,k}^{2}\cdot b_{k^{\prime},j^{\prime}}^{2}$
$=\frac{\rho_{n}(G)-n^{3}}{n^{4}}\left\|\mathbf{A}\right\|_{F}^{2}\cdot\left\|\mathbf{B}\right\|_{F}^{2}$
where we used the fact that the number of non-trivial collisions is $\rho_{n}(G)-n^{3}$ by hypothesis, and thus the probability of a quad-tuple $(i,j,k,k^{\prime})$ being an error term is $\frac{\rho_{n}(G)-n^{3}}{n^{4}}$ (because of the random shifts, this probability is uniform). This finishes the proof. ∎
Appendix D Fourier Domain Concentration Proofs
In this section we provide the proofs we omitted when proving Theorem 9.6.
Proof of 9.7.
The proof is symmetric in the case of $b$, so without loss of generality we focus on $a$. By definition of the Fourier transform (and using the fact $a$ takes real values):
$\mathbb{E}[\hat{a}(\xi)\cdot\overline{\hat{a}(\eta)}]$ $=\mathbb{E}\left[\left(\frac{1}{m^{3/2}}\sum_{x\in\mathrm{supp}(a)}a(x)\cdot\omega^{-x\cdot\xi}\right)\cdot\left(\frac{1}{m^{3/2}}\sum_{y\in\mathrm{supp}(a)}a(y)\cdot\omega^{y\cdot\eta}\right)\right]$
$=\frac{1}{m^{3}}\sum_{x,y\in\mathrm{supp}(a)}\mathbb{E}[a(x)\cdot a(y)]\cdot\omega^{-x\cdot\xi+y\cdot\eta}$
$=\frac{1}{m^{3}}\sum_{x\in\mathrm{supp}(a)}\underbrace{\mathbb{E}[a(x)^{2}]}_{\nu_{A}}\cdot\omega^{x\cdot(\eta-\xi)}+\frac{1}{m^{3}}\sum_{x\not=y\in\mathrm{supp}(a)}\underbrace{\mathbb{E}[a(x)]\cdot\mathbb{E}[a(y)]}_{\mu_{A}^{2}}\cdot\omega^{-x\cdot\xi+y\cdot\eta}$
$=\frac{1}{m^{3}}\cdot(\nu_{A}-\mu_{A}^{2})\cdot\sum_{x\in\mathrm{supp}(a)}\omega^{x\cdot(\eta-\xi)}+\frac{1}{m^{3}}\cdot\mu_{A}^{2}\cdot\sum_{x,y\in\mathrm{supp}(a)}\omega^{-x\cdot\xi+y\cdot\eta}.$
Let us examine the first sum of roots of unity. Denote $\delta=\xi-\eta$ and let $\left\lVert\delta\right\rVert_{0}$ denote the Hamming weight (number of non-zero coordinates). By 9.5, we can easily compute the sum:
$\sum_{x\in\mathrm{supp}(a)}\omega^{x\cdot(\eta-\xi)}=m^{3/2}\cdot\hat{f}_{a}(\delta)=\begin{cases}2(m-1)^{2}&\delta=0,\\ (m-1)^{2}-(m-1)&\delta_{1}=\delta_{2}=0\lor\delta_{2}=\delta_{3}=0,\\ -2(m-1)&\delta_{1}=\delta_{3}=0,\\ -(m-1)+1&\delta_{1}=0\lor\delta_{3}=0,\\ 2&\text{otherwise}.\end{cases}$
Similarly, we compute the second sum:
$\sum_{x,y\in\mathrm{supp}(a)}\omega^{-x\cdot\xi+y\cdot\eta}=\left(\sum_{x}\omega^{-x\cdot\xi}\right)\cdot\left(\sum_{y}\omega^{y\cdot\eta}\right)=m^{3}\cdot\hat{f}_{a}(\xi)\cdot\hat{f}_{a}(-\eta).$
Since $\xi,\eta\notin K$ we know that $\hat{f}_{a}(\xi)=\hat{f}_{a}(-\eta)=\frac{2}{m^{3/2}}$ (note that $\eta\notin K$ implies $\eta$ is not contained in any of the heavy planes or axes), hence $m^{3}\cdot\hat{f}_{a}(\xi)\cdot\hat{f}_{a}(-\eta)=4.$ We conclude that
$\mathbb{E}[\hat{a}(\xi)\cdot\overline{\hat{a}(\eta)}]=\frac{4\mu_{A}^{2}}{m^{3}}+\underbrace{(\nu_{A}-\mu_{A}^{2})}_{=\sigma_{A}^{2}}\times\begin{cases}\frac{2(m-1)^{2}}{m^{3}}&\delta=0,\\ \frac{(m-1)^{2}-(m-1)}{m^{3}}&\delta_{1}=\delta_{2}=0\lor\delta_{2}=\delta_{3}=0,\\ \frac{-2(m-1)}{m^{3}}&\delta_{1}=\delta_{3}=0\\ \frac{-m+2}{m^{3}}&\delta_{1}=0\lor\delta_{3}=0,\\ \frac{2}{m^{3}}&\text{otherwise}.\end{cases}$
This shows $\mathbb{E}[\hat{a}(\xi)\cdot\overline{\hat{a}(\eta)}]$ is just a function of $\delta$, which we denote by $C_{a}$. The estimates now follow directly. ∎
Proof of 9.8.
We split to cases:
  • $\left\lVert\delta\right\rVert_{0}=0$: Then
    $T(\delta)=\sum_{0\not=x,y\in\mathbb{Z}_{m}}\omega^{0}+\sum_{x,z\in\mathbb{Z}_{m}}\omega^{0}=2(m-1)^{2}.$
  • $\left\lVert\delta\right\rVert_{0}=1$:
    • Suppose $\delta_{1}\not=0$. Then
      $T(\delta)=\sum_{0\not=x,y\in\mathbb{Z}_{m}}\omega^{x\cdot\delta_{1}}+\sum_{x,z\in\mathbb{Z}_{m}}\omega^{z\cdot\delta_{1}}=2(-1)(m-1)=2-2m.$
    • Suppose $\delta_{2}\not=0$. Then
      $T(\delta)=\sum_{0\not=x,y\in\mathbb{Z}_{m}}\omega^{y\cdot\delta_{2}}+\sum_{x,z\in\mathbb{Z}_{m}}\omega^{0}=(m-1)^{2}-(m-1).$
    • Suppose $\delta_{3}\not=0$. Then
      $T(\delta)=\sum_{0\not=x,y\in\mathbb{Z}_{m}}\omega^{0}+\sum_{x,z\in\mathbb{Z}_{m}}\omega^{z\cdot\delta_{3}}=(m-1)^{2}-(m-1).$
  • $\left\lVert\delta\right\rVert_{0}=2$:
    • Suppose $\delta_{1},\delta_{2}\not=0,\delta_{3}=0$. Then
      $T(\delta)=\sum_{0\not=x,y\in\mathbb{Z}_{m}}\omega^{x\cdot\delta_{1}+y\cdot\delta_{2}}+\sum_{0\not=x,z\in\mathbb{Z}_{m}}\omega^{x\cdot\delta_{1}}=(-1)^{2}+(m-1)\cdot(-1)=-m+2.$
    • Suppose $\delta_{1},\delta_{3}\not=0,\delta_{2}=0$. Then
      $T(\delta)=\sum_{0\not=x,y\in\mathbb{Z}_{m}}\omega^{x\cdot\delta_{1}}+\sum_{0\not=x,z\in\mathbb{Z}_{m}}\omega^{x\cdot\delta_{1}+z\cdot\delta_{3}}=-m+2.$
    • Suppose $\delta_{2},\delta_{3}\not=0,\delta_{1}=0$. Then
      $T(\delta)=\sum_{0\not=x,y\in\mathbb{Z}_{m}}\omega^{y\cdot\delta_{2}}+\sum_{0\not=x,z\in\mathbb{Z}_{m}^{3}}\omega^{z\cdot\delta_{3}}=2(m-1)\cdot(-1)=2-2m.$
  • $\left\lVert\delta\right\rVert_{0}=3$: It holds
    $T(\delta)=\sum_{0\not=x,y\in\mathbb{Z}_{m}}\omega^{x\cdot\delta_{1}+y\cdot\delta_{2}}+\sum_{0\not=x,z\in\mathbb{Z}_{m}}\omega^{x\cdot\delta_{1}+z\cdot\delta_{3}}=2(-1)^{2}=2.$
The estimates follow directly from these computations. ∎
Proof of 9.10.
Let $Y=\left\lVert a\right\rVert_{2}^{2}\cdot\left\lVert b\right\rVert_{2}^{2}$ and denote $W=\left\lVert a\right\rVert_{2}^{2},Z=\left\lVert b\right\rVert_{2}^{2}$. Since $W,Z$ are independent random variables, we have $\mathbb{E}[Y]=\mathbb{E}[W]\cdot\mathbb{E}[Z]$. By definition,
$\mathbb{E}[W]=\sum_{x\in\mathrm{supp}(a)}\mathbb{E}[a_{x}^{2}]=\left|\mathrm{supp}(a)\right|\cdot\nu_{A}=2(m-1)^{2}\cdot\nu_{A},$
and similarly $\mathbb{E}[Z]=2(m-1)^{2}\cdot\nu_{B}$. Thus
$\mathbb{E}[Y]=4(m-1)^{4}\cdot\nu_{A}\nu_{B}.$
Moreover, since $W,Z$ are both sums of independent random variables, with variance proxies $p_{A}^{2},p_{B}^{2}$, we obtain that $W,Z$ have variance proxies $2(m-1)^{2}\cdot p_{A}^{2}$ and $2(m-1)^{2}\cdot p_{B}^{2}$ respectively. Therefore, by Hoeffding’s inequality (stated as: for i.i.d. $X_{1},\ldots,X_{n}$ sub-Gaussians with variance proxy $p^{2}$, it holds $\forall\varepsilon\geq 0$, $\Pr\left(\left|\sum_{i}(X_{i}-\mathbb{E}[X_{i}])\right|\geq\varepsilon\right)\leq 2e^{-\varepsilon^{2}/(2n\cdot p^{2})}$):
$\Pr(W<\delta\mathbb{E}[W])$ $\leq\Pr(|W-\mathbb{E}[W]|\geq(1-\delta)\mathbb{E}[W])$
$\leq 2\exp\left(-\frac{(1-\delta)^{2}\mathbb{E}[W]^{2}}{4\cdot(m-1)^{2}\cdot p_{A}^{2}}\right)$
$=2\exp\left(-\frac{(1-\delta)^{2}\cdot\nu_{A}^{2}\cdot(m-1)^{2}}{p_{A}^{2}}\right).$
A similar bound holds for $Z$, using $B$-parameters instead. Since $Y=W\cdot Z$, we obtain that $Y<\varepsilon\cdot\mathbb{E}[Y]$ implies $W<\sqrt{\varepsilon}\cdot\mathbb{E}[W]$ or $Z<\sqrt{\varepsilon}\cdot\mathbb{E}[Z]$. Hence
$\Pr(Y<\varepsilon\mathbb{E}[Y])$ $\leq\Pr(W<\sqrt{\varepsilon}\mathbb{E}[W])+\Pr(Z<\sqrt{\varepsilon}\mathbb{E}[Z])$
$\leq 2\exp\left(-\frac{(1-\delta)^{2}\cdot\nu_{A}^{2}\cdot(m-1)^{2}}{p_{A}^{2}}\right)+2\exp\left(-\frac{(1-\delta)^{2}\cdot\nu_{B}^{2}\cdot(m-1)^{2}}{p_{B}^{2}}\right).$
Set $\varepsilon=\frac{1}{2}$, then $(1-\sqrt{\varepsilon})^{2}=\frac{3}{2}-\sqrt{2}$. Therefore setting $\alpha=(6-4\sqrt{2})\approx 0.34$ we have
$\Pr\left(Y<\frac{1}{2}\mathbb{E}[Y]\right)\leq 2\exp\left(-(m-1)^{2}\cdot\alpha\cdot(\nu_{A}/p_{A})^{2}\right)+2\exp\left(-(m-1)^{2}\cdot\alpha\cdot(\nu_{B}/p_{B})^{2}\right)\leq 4\exp(-\gamma\cdot(m-1)^{2}),$
where $\gamma=\alpha\cdot\min\left(\frac{\nu_{A}^{2}}{p_{A}^{2}},\frac{\nu_{B}^{2}}{p_{B}^{2}}\right)$. ∎
Lemma D.1.
Suppose we have an algorithm that given matrices $\mathbf{A}_{1},\mathbf{A}_{2},\mathbf{B}_{1},\mathbf{B}_{2}\in\mathbb{R}^{d\times d}$ computes an approximation $\mathbf{C}_{1},\mathbf{C}_{2}$ such that
$\left\lVert\mathbf{A}_{1}\mathbf{B}_{1}-\mathbf{C}_{1}\right\rVert_{2}^{2}+\left\lVert\mathbf{A}_{2}\mathbf{B}_{2}-\mathbf{C}_{2}\right\rVert_{2}^{2}\leq\varepsilon\cdot\left(\left\lVert\mathbf{A}_{1}\right\rVert_{2}^{2}+\left\lVert\mathbf{A}_{2}\right\rVert_{2}^{2}\right)\cdot\left(\left\lVert\mathbf{B}_{1}\right\rVert_{2}^{2}+\left\lVert\mathbf{B}_{2}\right\rVert_{2}^{2}\right).$
Then approximating the product of two matrices $\mathbf{A},\mathbf{B}\in\mathbb{R}^{2d\times 2d}$ can be done with error $\left\lVert\mathbf{A}\mathbf{B}-\mathbf{C}\right\rVert_{F}^{2}\leq 2\varepsilon\cdot\left\lVert\mathbf{A}\right\rVert_{F}^{2}\cdot\left\lVert\mathbf{B}\right\rVert_{F}^{2}$.
Proof.
Partition each matrix into $4$ blocks, of size $d\times d$, writing
$\mathbf{A}=\left(\begin{smallmatrix}\mathbf{A}_{1,1}&\mathbf{A}_{1,2}\\ \mathbf{A}_{2,1}&\mathbf{A}_{2,2}\end{smallmatrix}\right)\quad,\quad\mathbf{B}=\left(\begin{smallmatrix}\mathbf{B}_{1,1}&\mathbf{B}_{1,2}\\ \mathbf{B}_{2,1}&\mathbf{B}_{2,2}\end{smallmatrix}\right).$
Then $\mathbf{A}\mathbf{B}$ can be written as
$\left(\begin{smallmatrix}\left(\mathbf{A}_{1,1}\mathbf{B}_{1,1}+\mathbf{A}_{1,2}\mathbf{B}_{2,1}\right)&\left(\mathbf{A}_{1,1}\mathbf{B}_{1,2}+\mathbf{A}_{1,2}\mathbf{B}_{2,2}\right)\\ \left(\mathbf{A}_{2,1}\mathbf{B}_{1,1}+\mathbf{A}_{2,2}\mathbf{B}_{2,1}\right)&\left(\mathbf{A}_{2,1}\mathbf{B}_{1,2}+\mathbf{A}_{2,2}\mathbf{B}_{2,2}\right)\end{smallmatrix}\right).$
For each $i,j\in\left\{1,2\right\}$, apply the given algorithm on $\mathbf{A}_{i,1},\mathbf{A}_{i,2},\mathbf{B}_{1,j},\mathbf{B}_{2,j}$ to obtain two approximate products $\mathbf{D}^{1}_{i,j},\mathbf{D}^{2}_{i,j}$ and set the $(i,j)$-block of the result matrix $\mathbf{C}$ to be $\mathbf{C}_{i,j}=\mathbf{D}_{1}+\mathbf{D}_{2}$. Since the blocks are disjoint, we have
$\left\lVert\mathbf{C}-\mathbf{A}\mathbf{B}\right\rVert_{F}^{2}=\sum_{i,j\in\left\{1,2\right\}}\left\lVert\mathbf{C}_{i,j}-(\mathbf{A}\mathbf{B})_{i,j}\right\rVert_{F}^{2},$
and by the parallelogram identity, it holds
$\left\lVert\mathbf{C}_{i,j}-(\mathbf{A}\mathbf{B})_{i,j}\right\rVert_{F}^{2}=\left\lVert(\mathbf{D}_{i,j}^{1}-\mathbf{A}_{i,1}\mathbf{B}_{1,j})+(\mathbf{D}_{i,j}^{2}-\mathbf{A}_{i,2}\mathbf{B}_{2,j})\right\rVert_{F}^{2}\leq 2\left\lVert\mathbf{D}_{i,j}^{1}-\mathbf{A}_{i,1}\mathbf{B}_{1,j}\right\rVert_{F}^{2}+2\left\lVert\mathbf{D}^{2}_{i,j}-\mathbf{A}_{i,2}\mathbf{B}_{2,j}\right\rVert_{F}^{2}.$
Applying the algorithm’s error guarantee we obtain
$\left\lVert\mathbf{C}-\mathbf{A}\mathbf{B}\right\rVert_{F}^{2}$ $\leq 2\sum_{i,j\in\left\{1,2\right\}}\left(\left\lVert\mathbf{D}_{i,j}^{1}-\mathbf{A}_{i,1}\mathbf{B}_{1,j}\right\rVert_{F}^{2}+\left\lVert\mathbf{D}^{2}_{i,j}-\mathbf{A}_{i,2}\mathbf{B}_{2,j}\right\rVert_{F}^{2}\right)$
$\leq 2\varepsilon\sum_{i,j\in\left\{1,2\right\}}\left(\left\lVert\mathbf{A}_{i,1}\right\rVert_{F}^{2}+\left\lVert\mathbf{A}_{i,2}\right\rVert_{F}^{2}\right)\cdot\left(\left\lVert\mathbf{B}_{1,j}\right\rVert_{F}^{2}+\left\lVert\mathbf{B}_{2,j}\right\rVert_{F}^{2}\right)$
$=2\varepsilon\underbrace{\left(\sum_{i\in\left\{1,2\right\}}\left(\left\lVert\mathbf{A}_{i,1}\right\rVert_{F}^{2}+\left\lVert\mathbf{A}_{i,2}\right\rVert_{F}^{2}\right)\right)}_{=\left\lVert\mathbf{A}\right\rVert_{F}^{2}}\cdot\underbrace{\left(\sum_{j\in\left\{1,2\right\}}\left(\left\lVert\mathbf{B}_{1,j}\right\rVert_{F}^{2}+\left\lVert\mathbf{B}_{2,j}\right\rVert_{F}^{2}\right)\right)}_{=\left\lVert\mathbf{B}\right\rVert_{F}^{2}}$
$=2\varepsilon\cdot\left\lVert\mathbf{A}\right\rVert_{F}^{2}\cdot\left\lVert\mathbf{B}\right\rVert_{F}^{2}.$
Proof of 9.12.
  1. 1.
    Here we use the specific structure of $a$ and $b$. As we’ve discussed in item 1, the Fourier transforms of the partial signals $a_{1},a_{2},b_{1},b_{2}$ are fixed along a complementary axis. Let us focus on $a$, as the argument for $b$ is symmetric. Then $\hat{a}=\hat{a}_{1}+\hat{a}_{2}$ by linearity, and $\hat{a}_{1}$ is fixed on the $3$-rd axis ($z$-axis), while $\hat{a}_{2}$ is fixed on the $1$-st axis ($x$-axis). Computing the values of $\hat{a}_{1}$ on the $x,y$-plane can be done by a single $2$-dimensional FFT of size $m\times m$, with proper scaling. Similarly, computing $\hat{a}_{2}$ on the $y,z$-plane can be done by a single $2$-dimensional FFT of size $m\times m$. Both have runtime of $O(m^{2}\log m)$. Now, computing the frequencies in $K$ is just a matter of accessing these $m\times m$ arrays and adding two terms, so it can be done in $O(rm^{2})=O(\left|K\right|)$ time.
  2. 2.
    Recall that $K=K_{1}\cup K_{2}\cup K_{3}$, where $K_{i}=\left\{\xi\in\mathbb{Z}_{m}^{3}\mid\xi_{i}<r\right\}$. Let us focus on $K_{1}$, as the argument for $K_{2},K_{3}$ is the same, and if we show how to compute $\mathcal{F}^{-1}(\hat{f}|_{K_{i}})|_{U}$ then by linearity we can compute $\mathcal{F}^{-1}(\hat{f}|_{K})|_{U}$ by summing the results in a specific way. It is convenient to think of $K_{1}$ as a slab (as depicted in Figure 2), consisting of $r$ copies of the $y,z$-plane. Fix $\xi_{1}$, and let $\mathsf{YZ}_{\xi_{1}}:=(\xi_{1},*,*)$ denote the $y,z$ plane corresponding to $\xi_{1}$. Then for $u\in U$:
    $\mathcal{F}^{-1}(\hat{f}|_{\mathsf{YZ}_{\xi_{1}}})_{u}=\frac{1}{m^{3/2}}\sum_{\xi_{2},\xi_{3}\in[m]}\hat{f}(\xi_{1},\xi_{2},\xi_{3})\cdot\omega^{\xi_{1}u_{1}+\xi_{2}u_{2}+\xi_{3}u_{3}}=\frac{\omega^{\xi_{1}u_{1}}}{\sqrt{m}}\mathcal{F}^{-1}_{\mathbb{Z}_{m}^{2}}(\hat{f}_{\xi_{1}})_{(u_{2},u_{3})},$
    where $\hat{f}_{\xi_{1}}:\mathbb{Z}_{m}^{2}\to\mathbb{C}$ is the function $\hat{f}_{\xi_{1}}(\xi_{2},\xi_{3})=\hat{f}(\xi_{1},\xi_{2},\xi_{3})$. In other words, computing the Fourier transform of $\hat{f}|_{\mathsf{YZ}_{\xi_{1}}}$ reduces to computing a $2$-dimensional FFT of size $m\times m$. Suppose we compute these FFTs for every $0\leq\xi_{1}<r$, then using the fact $K_{1}=\bigsqcup_{\xi_{1}<r}\mathsf{YZ}_{\xi_{1}}$, we obtain that
    $\mathcal{F}^{-1}(\hat{f}|_{K_{1}})_{u}=\sum_{\xi_{1}<r}\mathcal{F}^{-1}(\hat{f}|_{\mathsf{YZ}_{\xi_{1}}})_{u},$
    by linearity. Thus, to compute $\mathcal{F}^{-1}(\hat{f}|_{K_{1}})$ we need to execute $r$ FFTs of size $m\times m$, each costing $O(m^{2}\log m)$. Assuming we can access these lower dimensional FFTs, in order to compute the coefficient of a specific element $u\in U$, we need to lookup $r$ values, multiply them by weights, and sum the result. Therefore, computing all elements in $U$ can be done in $\left|U\right|\cdot r$ time. The total runtime is therefore established.